### v1

# Supersymmetric BMS4 Algebras Revisited: Electric/Magnetic Superalgebras and Free Field Realization

Abstract
In this work, we present a systematic classification of supersymmetric extensions of the BMS4 algebra and their realizations in free field theories. By requiring that supercharges admit finite-dimensional subsectors, we identify ten distinct electric super BMS4 algebras and six magnetic ones. The electric case is characterized by supercharge anticommutators closing on supertranslations, while the magnetic case necessarily involves superrotations. To realize these algebras in free field theories, we follow a constructive procedure: first identify the modes bilinears of which generate the symmetry algebra, then determine the fields with appropriate transformation properties under the BMS4 algebra, and finally construct consistent theories whose equations of motion admit the desired supersymmetry. Notably, $R$-symmetry with nonvanishing spin is essential for the Type II-II, Type I-II, and Type II-I magnetic super BMS4 algebras, shedding new light on spacetime structure of string theory for flat holography. Moreover, in the Type I-I theory, $R$-symmetry relates electric and magnetic scalars, indicating their equal significance. In addition, $R$-symmetry maps the electric scalar to a spin-$1$ field and vice versa, offering a novel perspective on supersymmetric extensions of soft theorems.
a Beijing Institute of Mathematical Sciences and Applications (BIMSA), Huaibei Town, Huairou District, Beijing 101408, China
1 Introduction
The study of asymptotic symmetries in general relativity originates from the seminal work of Bondi, van der Burg, Metzner, and Sachs [1, 2, 3, 4]. They demonstrated that the symmetry group of 44D asymptotically flat spacetime at null infinity extends beyond the Poincaré group. This enlarged symmetry is now referred to as the BMS4 algebra. A central feature of this algebra is the infinite-dimensional family of generators known as supertranslations, which extend the usual translations in Minkowski space. Beyond supertranslations, additional structure was subsequently identified. A systematic formulation of these “superrotations” was developed by Barnich and Troessaert [5, 6, 7, 8], who showed that they can be interpreted as local conformal transformations of the celestial sphere. In this sense, superrotations generalize global Lorentz transformations in the same way that supertranslations generalize translations. The resulting structure, often called the extended BMS algebra, takes the form
$[L_{m},L_{n}]=(m-n)L_{m+n},\quad[\bar{L}_{\bar{m}},\bar{L}_{\bar{n}}]=(\bar{m}-\bar{n})\bar{L}_{\bar{m}+\bar{n}},$ (1.1)
$[L_{m},M_{r,\bar{r}}]=\left(\frac{m}{2}-r\right)M_{m+r,\bar{r}},\quad[\bar{L}_{\bar{m}},M_{r,\bar{r}}]=\left(\frac{\bar{m}}{2}-\bar{r}\right)M_{r,\bar{m}+\bar{r}},$
where the superrotations $L,\bar{L}$ and the supertranslations $M$ together generate an infinite-dimensional symmetry. This extension has significant implications for gravitational scattering, and later work established its relation to subleading soft graviton theorems [9, 10]. For recent reviews, see [11, 12, 13].
For several decades the BMS algebra remained primarily a mathematical curiosity, until renewed interest arose from its role in infrared physics. In particular, Strominger and collaborators demonstrated that the Ward identities associated with BMS supertranslations reproduce Weinberg’s soft graviton theorem [14, 15, 16, 17, 18, 19, 20]. This correspondence established a precise relation among asymptotic symmetries, soft theorems, and memory effects, often referred to as the infrared triangle [13]. In recent years, BMS4 symmetries have become central to the framework of celestial holography. Within this approach, scattering amplitudes in asymptotically flat spacetime are reinterpreted as correlation functions of conformal primaries defined on the celestial sphere [21, 22, 23, 24, 25, 26, 27]. In this setting, the BMS4 algebra provides the organizing principle of the theory, much like the Virasoro and affine Kac–Moody algebras in 22D conformal field theory. Current research continues to develop its representation theory [28, 29, 30, 31] and investigate its significance for flat holography [7, 22, 32, 27, 33], quantum gravity [14, 9], and even black hole physics [34, 35, 36].
While the BMS4 algebra encodes the asymptotic symmetries of flat spacetime, it is natural to ask whether these structures admit supersymmetric extensions. Supersymmetry plays a central role in modern theoretical physics, appearing in supergravity, string theory, and proposals for ultraviolet-complete quantum gravity theories. Extending BMS4 by incorporating fermionic generators enables the study of asymptotic symmetries in supergravity and the analysis of constraints imposed by supersymmetry at asymptotic infinities [37, 38, 39, 40]. Moreover, such extensions provide a systematic framework for examining celestial holography in supersymmetric regimes, offering insights into scattering amplitudes involving gravitinos and their superpartners [23, 41, 42, 43, 44]. The simplest extension introduces spin-$\frac{1}{2}$ generators $Q_{\alpha}$ alongside the bosonic supertranslation and superrotation generators [37, 45, 46, 47, 38, 48, 49, 50]. These generators satisfy anticommutation relations consistent with the Poincaré supersymmetry algebra, while commuting appropriately with the extended BMS generators. Schematically, the supersymmetric BMS algebra contains generators $\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{r,\bar{r}},Q^{(0,\frac{1}{2})}_{r,\bar{r}}\}$222 $(l,\bar{l})$ for the super charge label the $\mathfrak{so}(3)$ spins in Lorentzian group $\mathfrak{so}(3)\times\mathfrak{so}(3)$. See Section 2.1 for more details. , and the commutation relation besides (1.1) is
$\{Q^{(\frac{1}{2},0)}_{r,\bar{r}},Q^{(0,\frac{1}{2})}_{s,\bar{s}}\}=M_{r+s,\bar{r}+\bar{s}}.$ (1.2)
This algebra illustrates how supersymmetry intertwines with the infinite-dimensional asymptotic symmetries, thereby generalizing the infrared structures identified in purely bosonic theories.
However, the above is not the full story of supersymmetric extension of the BMS4 algebra. In this work, we identify additional supersymmetric extensions, which can be classified as electric and magnetic types. Specifically, the electric super BMS4 algebra refers to the case where anticommutators of supercharges produce only supertranslations, while the magnetic super BMS4 algebra is characterized by anticommutators that also yield superrotations (See [51, 52] for early discussion on magnetic examples). In this classification, the simplest extension (1.2) corresponds to an electric super BMS4 algebra. This paper focuses on solutions admitting finite-dimensional subalgebras, which have potential applications in flat holography [53, 54, 23, 55]. For the electric case, there are ten distinct unextended superalgebras. It was realized in section 3 that only three of them can be realized without introducing spin higher than $\frac{1}{2}$ in the corresponding symmetry representations. The magnetic cases are more intricate. There are two types of chiral super BMS4 algebras: Type I (2.9ac) and Type II (2.9ad), along with four nonchiral superalgebras formed by combinations of these two types. Notably, for the nonchiral cases, $R$-symmetry is necessary for the closure of the algebra. The representation of the $R$-charges varies in each case, depending on the types of left- and right-handed supercharges. The appearance of the $R$-charges indicates that these combined super BMS4 algebras should be interpreted as extended supersymmetry.
Beyond the algebraic formulation, this study presents an explicit free field realization of the super BMS4 algebra, by expressing the algebra generators as bilinears of the field modes. Specifically, we construct the supersymmetric extension of the massless electric scalar theory $\mathcal{L}=\frac{1}{2}\partial_{0}\phi\partial_{0}\phi$, representing the simplest scalar multiplet. For the electric superalgebras, the theory (2.9y) admits three distinct electric superalgebras; conversely, for the magnetic superalgebras, each constructed theory is unique. Notably, although theory (2.9be) resembles the electric theory (2.9y), they are fundamentally different due to the distinct conformal dimensions of the fermionic fields. Interestingly, theory (2.9ca) establishes a relationship between the electric and magnetic scalars via $R$-symmetry, highlighting the essential role of magnetic scalars in supersymmetric flat holography. In theory (2.9dy), the emergence of a spin-$1$ massless boson suggests potential applications in supersymmetric soft theorems. Moreover, part of the theory (2.9dy) appeared in [21] as free perturbative supergravity theory. These realizations establish a direct correspondence between the algebraic structure of super BMS4 and physical observables, thereby facilitating the exploration of supersymmetric celestial conformal field theories [56] and the investigation of how supersymmetry influences the infrared behavior of gravitational scattering processes. Such theories serve as fundamental building blocks for understanding the spectrum and dynamics of supersymmetric theories in asymptotically flat spacetime. Moreover, the free field approach provides a systematic framework for deriving operator product expansions and correlation functions, which are essential for studying holographic dualities and soft theorems in supersymmetric gravity.
The remainder of this paper is organized as follows. In Section 2, we discuss the possible super BMS4 algebras. After introducing the charge representations in Section 2.1, we identify ten electric super BMS4 algebras in Section 2.2, along with two chiral magnetic super BMS4 algebras and four nonchiral superalgebras in Section 2.3. We then proceed to the free field realizations of these algebras in Section 3. Specifically, in Section 3.2, we demonstrate that theory (2.9y) is trivial in the sense that it preserves three of the electric superalgebras. In the magnetic sector discussed in Section 3.3, we show that the field theories admitting different superalgebras are distinct. Lastly, we conclude this work in Section 4. Appendix A contains detailed calculations of the linear difference equations arising from the constraints on the representations.
2 Super BMS4 Algebras
It is well known that the global part of the BMS4 algebra is the 33D Carrollian conformal algebra [57]. As established in the literature, every aspect of Carrollian theories, such as the limiting procedures and types of field theories, has electric and magnetic versions [58, 57, 59]. In this paper, we adopt the terminology from Carrollian theory to classify the super BMS4 algebras. Specifically, we define a super BMS4 algebra as electric if the anticommutators of the supercharges yield only supertranslations. In contrast, we classify a super BMS4 algebra as magnetic if the anticommutators produce both superrotations and supertranslations:
electric: $\{Q,Q\}\propto M$ (2.1)
magnetic: $\{Q,Q\}\propto L,\bar{L}\text{ or }M$
In this section, we systematically derive the super BMS4 algebras under the following assumptions. First, we assume that supersymmetry is generated by certain currents, implying that the supercharges reside in charge representations, which will be introduced in Section 2.1. Second, considering the potential applications in flat space holography, we require that the super BMS4 algebra possesses finite-dimensional subalgebras, which can be viewed as supersymmetric extensions of the Poincaré algebra. Lastly, to respect the spin-statistics theorem, we stipulate that the supercharges must have half-integer spins. With these assumptions, we classify the super BMS4 algebras into electric and magnetic types in Sections 2.2 and 2.3 respectively.
2.1 Charge representation of BMS4 algebra
In the context of supersymmetry, fermionic charges act as modes of supersymmetric currents. This subsection addresses the representation theory of these generators, which are associated with specific currents. For simplicity, we first examine the representation under a single branch of the Virasoro algebra and subsequently extend this to the full BMS4 algebra within this framework.
The symmetric generator $Q_{n}$ of a holomorphic current $J(z)$ with conformal dimension $h$ is expressed as
$J(z)=\sum_{r\in\mathbb{Z}+r_{0}}Q_{r}z^{-r-h}.$ (2.2)
The action of Virasoro algebra on $Q_{n}$ is given by
$L_{m},Q_{r}]=((h-1)m-r)Q_{m+r}.$ (2.3)
Verification of the Jacobi identity for $\{L_{m},L_{n},Q_{r}\}$ confirms that this defines a self consistent representation. In this work, we refer to this as the charge representation. Notably, the charge representation admits a finite subrepresentation if there exists a cutoff in the action of $L_{\pm 1}$ on $Q$:
$\exists~r>s\in\mathbb{Z}+r_{0}\text{ s.t. }[L_{1},Q_{r}]=[L_{-1},Q_{s}]=0.$ (2.4)
This amounts to solving (h1)r=(h1)s=0(h-1)-r=-(h-1)-s=0, which yields the following solutions:
[…]
For these cases, $\{Q_{h-1},\dots,Q_{-h+1}\}$ forms a finite-dimensional representation space of the subalgebra $\{L_{1},L_{0},L_{-1}\}=\mathfrak{so}(3)$ with spin-l=h1l=h-1. In the remainder of this paper, we use the notation $\mathfrak{so}(3)$ spin-$l$ (and $\bar{l}$ for the anti-holomorphic part) to label the charge representations, extending this notation to general $l\in\mathbb{R}$.
Considering a (super)algebra that contains multiple such symmetry charges $Q^{i}$, the Jacobi identities of the algebra should be saturated such that the algebra is well-defined. Assume that the commutation relation is
$Q^{i},Q^{j}\}\propto Q^{k},\quad[Q^{i},Q^{j}\}=\begin{cases}Q^{i}Q^{j}+Q^{j}Q^{i}&\text{$Q^{i}$ and $Q^{j}$ are fermionic,}\\ Q^{i}Q^{j}-Q^{j}Q^{i}&\text{otherwise.}\\ \end{cases}$ (2.6)
The Jacobi identity of $\{L_{0},Q^{i},Q^{j}\}$ requires that the index of $Q^{k}$ equals the sum of the indices of $Q^{i}$ and $Q^{j}$:
$Q^{i}_{r_{i}},Q^{j}_{r_{j}}\}=f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})~Q^{k}_{r_{i}+r_{j}},$ (2.7)
then the Jacobi identity of $\{L_{n},Q^{i},Q^{j}\}$ for generic $n$ further gives a constraint on the function $f$:
$(l_{k}n-(r_{i}+r_{j}))f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})-(l_{j}n-r_{j})f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j}+n)-(l_{i}n-r_{i})f_{l_{i},l_{j}}^{l_{k}}(r_{i}+n,r_{j})=0.$ (2.8)
This is a linear difference equation which is only defined on discrete $n,r_{i},r_{j}\in\mathbb{Z}$ points. Numerically, we find the following useful solutions333 Here we consider solutions with $l\in\mathbb{Z}/2$ and $l>0$, i.e., charge representations that admit finite-dimensional sub-representations. See Appendix A for more details.
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=1,$ $l_{k}=l_{i}+l_{j},$ (2.9a)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=l_{j}r_{i}-l_{i}r_{j},$ $l_{k}=l_{i}+l_{j}-1,$ (2.9b)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=r_{i}(l_{j}r_{i}+r_{j}),$
up to overall factors. Especially, (2.9b) applies to the BMS4 algebra itself. The case that $Q^{i}=Q^{j}=Q^{k}=L$, which corresponds to $l_{i}=l_{j}=1$ and $l_{k}=l_{i}+l_{j}-1=1$, gives the commutation relation
$L_{m},L_{n}]=(1*m-1*n)L_{m+n},$ (2.9j)
while the case that $Q^{i}=L,Q^{j}=Q^{k}=M$, which corresponds to $l_{i}=1,l_{j}=\frac{1}{2}$ and $l_{k}=l_{i}+l_{j}-1=\frac{1}{2}$, gives the commutation relation
$L_{m},M_{r,\bar{r}}]=\left(\frac{m}{2}-1*r\right)M_{m+r,\bar{r}}.$ (2.9k)
This result aligns with the fact that generators in the finite subalgebra $L_{m},\bar{L}_{\bar{m}},M_{r,\bar{r}}$ are respectively in $(1,0)$, $(0,1)$ and $\left(\frac{1}{2},\frac{1}{2}\right)$ representations under the Lorentzian algebra $\mathfrak{so}_{L}(3)\times\mathfrak{so}_{R}(3)$.
It is worth emphasizing that the tensor product rules in (2.9) deviate from the familiar decomposition of $\mathfrak{so}(3)$ spin representations. For the latter, one has the standard rule
$(l_{i})\otimes(l_{j})=(|l_{i}-l_{j}|)\oplus\cdots\oplus(l_{i}+l_{j}),$ (2.9l)
where all intermediate spins appear. In contrast, for the charge representations under consideration, not all such intermediate spins are admissible. Specifically, the condition $l_{k}<l_{i}+l_{j}-2$ fails to satisfy the Jacobi identity equation (2.8) for generic $l_{i}$ and $l_{j}$. This indicates that the charge representation enforces a more restrictive structure, reflecting the algebraic consistency required by supersymmetric extensions rather than the purely group theoretic rules of $\mathfrak{so}(3)$.
Building on the above discussion, a charge $Q$ in the BMS4 algebra can be characterized by its holomorphic and anti-holomorphic spins $(l,\bar{l})$, with commutation relations given by
$L_{m},Q^{(l,\bar{l})}_{r,\bar{r}}]=(lm-r)Q^{(l,\bar{l})}_{m+r,\bar{r}},\qquad[\bar{L}_{\bar{m}},Q^{(l,\bar{l})}_{r,\bar{r}}]=(\bar{l}\bar{m}-\bar{r})Q^{(l,\bar{l})}_{r,\bar{m}+\bar{r}}$. (2.9m)
When the supertranslation generators $M$ act nontrivially on $Q$, i.e. $M,Q^{a}]\propto Q^{b}\neq 0$, the spins of $M$ and $Q^{(i)}$ are constrained by the tensor product rule in (2.9). In this sense, a super BMS4 algebra can be systematically constructed by enlarging the BMS4 algebra with additional fermionic generators, provided that the extended commutation relations are consistent. The labels $(l,\bar{l})$ already determine the action of $L$ and $\bar{L}$ on these new generators, so the nontrivial part of the construction lies in specifying the commutators of $M$ with the fermionic charges, as well as the (anti-)commutators among the new generators themselves. Ensuring algebraic closure then requires verifying the Jacobi identities case by case. In the following subsections, we will present explicit constructions of the electric and magnetic super BMS4 algebras.
2.2 Electric super BMS4 algebra
In this subsection, we reformulate the electric super BMS4 algebra using the charge representation framework introduced in Section 2.1. A standard construction of the electric extension begins with the infinite-dimensional uplift of the super Poincaré algebra. The resulting extended algebra is generated by $\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{r},Q^{(0,\frac{1}{2})}_{\bar{r}}\}$, where $Q^{(\frac{1}{2},0)}_{r}$ and $Q^{(0,\frac{1}{2})}_{\bar{r}}$ denote the fermionic supercharges. Their defining relations are
$M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{s}]=[M_{r,\bar{r}},Q^{(0,\frac{1}{2})}_{\bar{s}}]=0,\qquad\{Q^{(\frac{1}{2},0)}_{r},Q^{(0,\frac{1}{2})}_{\bar{r}}\}=M_{r,\bar{r}}.$ (2.9n)
The largest finite subalgebra of this infinite extension is generated by
$\{L_{\pm 1,0},\bar{L}_{\pm 1,0},M_{\pm\frac{1}{2},\pm\frac{1}{2}},Q^{(\frac{1}{2},0)}_{\pm\frac{1}{2}},Q^{(0,\frac{1}{2})}_{\pm\frac{1}{2}}\}$, (2.9o)
which is isomorphic to the standard $\mathcal{N}=1$ super Poincaré algebra. This correspondence becomes explicit after relabeling the generators as
$J^{0,1}_{\text{P}}=\frac{i}{2}(-L_{1}-L_{-1}+\bar{L}_{1}+\bar{L}_{-1}),~J^{0,2}_{\text{P}}=\frac{1}{2}(L_{1}-L_{-1}+\bar{L}_{1}-\bar{L}_{-1}),~J^{0,3}_{\text{P}}=-L_{0}-\bar{L}_{0},$ (2.9p)
$J^{1,2}_{\text{P}}=-iL_{0}+i\bar{L}_{0},~J^{1,3}_{\text{P}}=\frac{i}{2}(-L_{1}+L_{-1}+\bar{L}_{1}-\bar{L}_{-1}),~J^{2,3}_{\text{P}}=\frac{1}{2}(L_{1}+L_{-1}+\bar{L}_{1}+\bar{L}_{-1}),$
$P^{0}_{\text{P}}=\frac{1}{\sqrt{2}}(M_{\frac{1}{2},\frac{1}{2}}+M_{-\frac{1}{2},-\frac{1}{2}}),\qquad P^{1}_{\text{P}}=\frac{i}{\sqrt{2}}(M_{\frac{1}{2},-\frac{1}{2}}-M_{-\frac{1}{2},\frac{1}{2}}),$
$P^{2}_{\text{P}}=\frac{1}{\sqrt{2}}(-M_{\frac{1}{2},-\frac{1}{2}}-M_{-\frac{1}{2},\frac{1}{2}}),\qquad P^{3}_{\text{P}}=\frac{1}{\sqrt{2}}(-M_{\frac{1}{2},\frac{1}{2}}+M_{-\frac{1}{2},-\frac{1}{2}}),$
$Q_{\frac{1}{2}}^{\text{P}}=i~2^{\frac{3}{4}}Q^{(0,\frac{1}{2})}_{-\frac{1}{2}},\quad Q_{-\frac{1}{2}}^{\text{P}}=2^{\frac{3}{4}}Q^{(0,\frac{1}{2})}_{\frac{1}{2}},\quad\bar{Q}_{\frac{1}{2}}^{\text{P}}=-i~2^{\frac{3}{4}}Q^{(\frac{1}{2},0)}_{-\frac{1}{2}},\quad\bar{Q}_{-\frac{1}{2}}^{\text{P}}=2^{\frac{3}{4}}Q^{(\frac{1}{2},0)}_{\frac{1}{2}}$,
With this identification, the set $\{J^{\mu\nu}_{\text{P}},P^{\mu}_{\text{P}},Q^{\text{P}}_{\alpha},\bar{Q}^{\text{P}}_{\dot{\alpha}}\}$ reproduces the standard super Poincaré algebra.
However, the algebra obtained in (2.9n) does not exhaust all possible constructions of electric super BMS4. To proceed more systematically, we adopt a general strategy based on the following guiding principles. First, the supercharges are required to transform in charge representations, ensuring consistency with the underlying current algebra structure. Second, the labels $(l,\bar{l})$ of the supercharges are taken to lie in $\mathbb{Z}/2$ with $l,\bar{l}\geq 0$, so that the resulting algebra admits a finite-dimensional subalgebra, which can be viewed as a supersymmetric extension of the Poincaré algebra. Finally, to respect the spin–statistics theorem, the total spin must satisfy $l+\bar{l}\in\mathbb{Z}+\frac{1}{2}$. These ansatzes provide a systematic framework to classify and construct possible electric super BMS4 algebras beyond the minimal realization.
The cases with a single set of supercharges fail to satisfy the ansatzes outlined above. In fact, there are precisely two possible constructions compatible with the constraint (2.9). The first is given by (2.9):
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{1}{4},\frac{1}{4})}_{r,\bar{r}}\},\text{ with }\{Q^{(\frac{1}{4},\frac{1}{4})}_{r,\bar{r}},Q^{(\frac{1}{4},\frac{1}{4})}_{s,\bar{s}}\}=M_{r+s,\bar{r}+\bar{s}},$ (2.9q)
which satisfies the condition (2.9a). The second possibility reads
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{3}{4},\frac{3}{4})}_{r,\bar{r}}\},\text{ with }\{Q^{(\frac{3}{4},\frac{3}{4})}_{r,\bar{r}},Q^{(\frac{3}{4},\frac{3}{4})}_{s,\bar{s}}\}=\left(\frac{3}{4}r-\frac{3}{4}s\right)\left(\frac{3}{4}\bar{r}-\frac{3}{4}\bar{s}\right)M_{r+s,\bar{r}+\bar{s}},$ (2.9r)
consistent with the constraint (2.9b). Although these constructions are algebraically consistent, neither of them admits a finite-dimensional subalgebra. As a result, they fall outside the class of super BMS4 algebras envisioned by our ansatzes and do not yield a physically viable supersymmetric extension.
We now turn to the case with two sets of supercharges. A representative example is provided by the extension of the super Poincaré algebra (2.9n), where both the holomorphic and anti-holomorphic labels satisfy the condition (2.9a). More generally, the supercharge $Q^{(\frac{1}{2},0)}_{r}$ can be extended to $Q^{(\frac{1}{2},0)}_{r,\bar{r}}$, with the original operator identified as $Q^{(\frac{1}{2},0)}_{r}=Q^{(\frac{1}{2},0)}_{r,0}$. This extension introduces additional super BMS4 charges while leaving the finite-dimensional subalgebra unchanged. In a completely analogous fashion, the supercharge $Q^{(0,\frac{1}{2})}_{\bar{r}}$ can be extended to $Q^{(0,\frac{1}{2})}_{r,\bar{r}}$. Following this logic, we obtain ten distinct realizations of the electric super BMS4 algebra, classified according to the rules (2.9), up to exchange of holomorphic and anti-holomorphic labels.
holomorphic (2.9a), anti-holomorphic (2.9a):
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{r,\bar{r}},Q^{(0,\frac{1}{2})}_{r,\bar{r}}\},\quad\{Q^{(\frac{1}{2},0)}_{r,\bar{r}},Q^{(0,\frac{1}{2})}_{s,\bar{s}}\}=M_{r+s,\bar{r}+\bar{s}}.$ (2.9s)
holomorphic (2.9a), anti-holomorphic (2.9b):
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{1}{2},1)}_{r,\bar{r}},Q^{(0,\frac{1}{2})}_{r,\bar{r}}\},\quad\{Q^{(\frac{1}{2},1)}_{r,\bar{r}},Q^{(0,\frac{1}{2})}_{s,\bar{s}}\}=\left(\frac{1}{2}\bar{r}-\bar{s}\right)M_{r+s,\bar{r}+\bar{s}},$ (2.9t)
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{r,\bar{r}},Q^{(0,\frac{3}{2})}_{r,\bar{r}}\},\quad\{Q^{(\frac{1}{2},0)}_{r,\bar{r}},Q^{(0,\frac{3}{2})}_{s,\bar{s}}\}=\frac{3}{2}\bar{r}M_{r+s,\bar{r}+\bar{s}}.$ (2.9u)
holomorphic (2.9a), anti-holomorphic (LABEL:eq:TensorProductOfVirReprsSolution3):
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{r,\bar{r}},Q^{(0,\frac{5}{2})}_{r,\bar{r}}\},\quad\{Q^{(\frac{1}{2},0)}_{r,\bar{r}},Q^{(0,\frac{5}{2})}_{s,\bar{s}}\}=\bar{r}\left(\frac{5}{2}\bar{r}+\bar{s}\right)M_{r+s,\bar{r}+\bar{s}}.$ (2.9v)
holomorphic (2.9b), anti-holomorphic (2.9b):
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{1}{2},1)}_{r,\bar{r}},Q^{(1,\frac{1}{2})}_{r,\bar{r}}\},\quad\{Q^{(\frac{1}{2},1)}_{r,\bar{r}},Q^{(1,\frac{1}{2})}_{s,\bar{s}}\}=\left(r-\frac{1}{2}s\right)\left(\frac{1}{2}\bar{r}-\bar{s}\right)M_{r+s,\bar{r}+\bar{s}},$ (2.9w)
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{r,\bar{r}},Q^{(1,\frac{3}{2})}_{r,\bar{r}}\},\quad\{Q^{(\frac{1}{2},0)}_{r,\bar{r}},Q^{(1,\frac{3}{2})}_{s,\bar{s}}\}=\left(r-\frac{1}{2}s\right)\frac{3}{2}\bar{r}M_{r+s,\bar{r}+\bar{s}},$ (2.9x)
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{3}{2},0)}_{r,\bar{r}},Q^{(0,\frac{3}{2})}_{r,\bar{r}}\},\quad\{Q^{(\frac{3}{2},0)}_{r,\bar{r}},Q^{(0,\frac{3}{2})}_{s,\bar{s}}\}=\frac{9}{4}r\bar{s}M_{r+s,\bar{r}+\bar{s}}.$ (2.9y)
holomorphic (2.9b), anti-holomorphic (LABEL:eq:TensorProductOfVirReprsSolution3):
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{r,\bar{r}},Q^{(1,\frac{5}{2})}_{r,\bar{r}}\},\quad\{Q^{(\frac{1}{2},0)}_{r,\bar{r}},Q^{(1,\frac{5}{2})}_{s,\bar{s}}\}=\left(r-\frac{1}{2}s\right)\bar{r}\left(\frac{5}{2}\bar{r}+\bar{s}\right)M_{r+s,\bar{r}+\bar{s}},$ (2.9z)
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{3}{2},0)}_{r,\bar{r}},Q^{(0,\frac{5}{2})}_{r,\bar{r}}\},\quad\{Q^{(\frac{3}{2},0)}_{r,\bar{r}},Q^{(0,\frac{5}{2})}_{s,\bar{s}}\}=\frac{3}{2}s\bar{r}\left(\frac{5}{2}\bar{r}+\bar{s}\right)M_{r+s,\bar{r}+\bar{s}}.$ (2.9aa)
holomorphic (LABEL:eq:TensorProductOfVirReprsSolution3), anti-holomorphic (LABEL:eq:TensorProductOfVirReprsSolution3):
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{5}{2},0)}_{r,\bar{r}},Q^{(0,\frac{5}{2})}_{r,\bar{r}}\},\quad\{Q^{(\frac{5}{2},0)}_{r,\bar{r}},Q^{(0,\frac{5}{2})}_{s,\bar{s}}\}=s\left(r+\frac{5}{2}s\right)\bar{r}\left(\frac{5}{2}\bar{r}+\bar{s}\right)M_{r+s,\bar{r}+\bar{s}}.$ (2.9ab)
All of the above algebras contain finite-dimensional subalgebras. However, except for the case (2.9s), the corresponding finite-dimensional supercharges necessarily include spin-$\frac{3}{2}$ or spin-$\frac{5}{2}$ components. Such higher-spin fermionic generators are not compatible with the Haag–Łopuszański–Sohnius theorem, which allows only spin-$\frac{1}{2}$ supercharges in a consistent extension of the Poincaré algebra. Therefore, the only physically admissible super Poincaré algebra is the finite-dimensional subalgebra of (2.9s).
2.3 Magnetic chiral super BMS4 algebras
In the magnetic case, it’s required the existence of supercharges $Q^{a}$ such that $\{Q^{a},Q^{b}\}\propto L$. Since $L$ is purely holomorphic, consistency demands that the supercharges anticommuting into $L$ must also be purely holomorphic. Otherwise, the relation $\{Q^{a}_{r,\bar{r}},Q^{b}_{s,\bar{s}}\}=G_{r+s,\bar{r}+\bar{s}}$ would generate unphysical bosonic generators $L_{n,\bar{n}}$. An analogous argument applies to the anti-holomorphic sector. Therefore, it suffices to consider $\{Q^{a}_{r},Q^{b}_{r}\}\propto L$ with $Q^{a}_{r}\in(l_{a},0)$ and $Q^{b}_{r}\in(l_{b},0)$, together with $\{Q^{c}_{\bar{r}},Q^{d}_{\bar{r}}\}\propto\bar{L}$ where $Q^{c}_{\bar{r}}\in(0,l_{c})$ and $Q^{d}_{\bar{r}}\in(0,l_{d})$.
Assuming a single set of supercharges $Q$ generates $L$, the only consistent choice is $Q\in\left(\frac{1}{2},0\right)$ such that $\{Q^{(\frac{1}{2},0)}_{r},Q^{(\frac{1}{2},0)}_{s}\}=L_{r+s}$. The Jacobi identity for $\{M,Q^{(\frac{1}{2},0)},Q^{(\frac{1}{2},0)}\}$ then implies a nonvanishing commutator $M,Q^{(\frac{1}{2},0)}]$. Applying the rules of (2.9), one finds two distinct chiral magnetic BMS4 algebras:
Type I $\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{r},Q^{(0,\frac{1}{2})}_{r,\bar{r}}\}$:
$M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{s}]=\left(\frac{1}{2}r-\frac{1}{2}s\right)Q^{(0,\frac{1}{2})}_{r+s,\bar{r}},\quad\{Q^{(\frac{1}{2},0)}_{r},Q^{(\frac{1}{2},0)}_{s}\}=L_{r+s},\quad\{Q^{(\frac{1}{2},0)}_{r},Q^{(0,\frac{1}{2})}_{s,\bar{s}}\}=M_{r+s,\bar{s}},$ (2.9ac)
Type II $\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{r},Q^{(1,\frac{1}{2})}_{r,\bar{r}}\}$:
$M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{s}]=Q^{(1,\frac{1}{2})}_{r+s,\bar{r}},\quad\{Q^{(\frac{1}{2},0)}_{r},Q^{(\frac{1}{2},0)}_{s}\}=L_{r+s},\quad\{Q^{(\frac{1}{2},0)}_{r},Q^{(1,\frac{1}{2})}_{s,\bar{s}}\}=-\left(r-\frac{1}{2}s\right)M_{r+s,\bar{s}},$ (2.9ad)
In Type I, $M,Q^{(\frac{1}{2},0)}]\propto Q^{(0,\frac{1}{2})}$ satisfies (2.9b), whereas in Type II, $M,Q^{(\frac{1}{2},0)}]\propto Q^{(1,\frac{1}{2})}$ satisfies (2.9a). Unlike the electric case, the supercharges $Q^{(\frac{1}{2},0)}_{r}$ in (2.9ac) and (2.9ad) are not extended to $Q^{(\frac{1}{2},0)}_{r,\bar{r}}$, thereby avoiding the introduction of nonphysical generators $L_{m,\bar{m}}$. In contrast, the holomorphic index $r$ of $Q^{(0,\frac{1}{2})}_{r,\bar{r}}$ in (2.9ac) is required for the nonvanishing commutator $M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{s}]=Q^{(1,\frac{1}{2})}_{r+s,\bar{r}}$.
However, a mixture of the above two types does not yield a physically consistent superalgebra. Specifically, if one assumes that $\{Q^{(\frac{1}{2},0)}_{r},Q^{(0,\frac{1}{2})}_{r,\bar{r}},Q^{(1,\frac{1}{2})}_{r,\bar{r}}\}$ are the supercharges, the Jacobi identities for $\{Q^{(\frac{1}{2},0)},Q^{(\frac{1}{2},0)},Q^{(0,\frac{1}{2})}_{r,\bar{r}}\}$ and for $\{Q^{(\frac{1}{2},0)},Q^{(\frac{1}{2},0)},Q^{(1,\frac{1}{2})}_{r,\bar{r}}\}$ become mutually contradictory, unless one introduces an additional bosonic generator $\tilde{M}^{(\frac{1}{2},\frac{1}{2})}$. The resulting extended algebra takes the form:
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},\tilde{M}_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{r},Q^{(0,\frac{1}{2})}_{r,\bar{r}},Q^{(1,\frac{1}{2})}_{r,\bar{r}}\}$, (2.9ae)
$[M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{s}]=\frac{a}{2}(r-s)Q^{(0,\frac{1}{2})}_{r+s,\bar{r}}+(1-a)Q^{(1,\frac{1}{2})}_{r+s,\bar{r}}$,
$[\tilde{M}_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{s}]=\frac{1-ab}{2c}(r-s)Q^{(0,\frac{1}{2})}_{r+s,\bar{r}}+\frac{(a-1)b}{c}Q^{(1,\frac{1}{2})}_{r+s,\bar{r}}$
$\{Q^{(\frac{1}{2},0)}_{r},Q^{(\frac{1}{2},0)}_{s}\}=L_{r+s},\quad\{Q^{(\frac{1}{2},0)}_{r},Q^{(0,\frac{1}{2})}_{s,\bar{s}}\}=bM_{r+s,\bar{s}}+c\tilde{M}_{r+s,\bar{s}},$
$\{Q^{(\frac{1}{2},0)}_{r},Q^{(1,\frac{1}{2})}_{s,\bar{s}}\}=\left(r-\frac{1}{2}s\right)\left(\frac{1-ab}{a-1}M_{r+s,\bar{s}}+\frac{ac}{a-1}\tilde{M}_{r+s,\bar{s}}\right).$
where $a,b,c$ are arbitrary constants. Importantly, the additional generator $\tilde{M}$ lacks a clear physical interpretation. Therefore, such a mixed construction does not correspond to a physically meaningful symmetry.
Observing (2.9ac), we note that $Q^{(\frac{1}{2},0)}$ and $Q^{(0,\frac{1}{2})}$ are not related by conjugation. This makes it clear that both (2.9ac) and (2.9ad) describe chiral superalgebras, as only the holomorphic Virasoro algebra $L_{n}$ appears in anticommutators of super charges. For this reason, (2.9ac) and (2.9ad) are considered as holomorphic in this paper. Anti-holomorphic counterparts $\{\bar{Q}^{(0,\frac{1}{2})}_{\bar{r}},\bar{Q}^{(\frac{1}{2},0)}_{r,\bar{r}}\}$ and $\{\bar{Q}^{(0,\frac{1}{2})}_{\bar{r}},\bar{Q}^{(\frac{1}{2},1)}_{r,\bar{r}}\}$ can be defined through conjugation, yielding super BMS4 algebra structures analogous to (2.9ac) and (2.9ad).
For a super BMS4 algebra containing both holomorphic and anti-holomorphic sectors, simply including the corresponding supercharges is insufficient. For instance, consider holomorphic supercharges $\{Q^{(\frac{1}{2},0)}_{r},Q^{(0,\frac{1}{2})}_{r,\bar{r}}\}$ and anti-holomorphic supercharges $\{\bar{Q}^{(0,\frac{1}{2})}_{\bar{r}},\bar{Q}^{(\frac{1}{2},0)}_{r,\bar{r}}\}$, both of Type I as in (2.9ac). The Jacobi identity for $\{Q^{(\frac{1}{2},0)},Q^{(\frac{1}{2},0)},\bar{Q}^{(\frac{1}{2},0)}\}$ is satisfied only if a bosonic generator $R^{(0,0)}$ is introduced, such that
$\{Q^{(\frac{1}{2},0)},\bar{Q}^{(\frac{1}{2},0)}\}\propto R^{(0,0)}.$ (2.9af)
Within the finite-dimensional subalgebra, $R=R^{(0,0)}_{0,0}$ naturally functions as an $R$-symmetry. Consequently, the super BMS4 algebra requires the presence of $R$-symmetry even when it is not an extended supersymmetry algebra. This motivates the construction of magnetic super BMS4 algebras combining different types of supercharges, and we label the resulting algebras Type A-B for the holomorphic part being Type A and anti-holomorphic part being Type B. The resulting superalgebras are:
Type I-I $\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{r},Q^{(0,\frac{1}{2})}_{r,\bar{r}},\bar{Q}^{(0,\frac{1}{2})}_{\bar{r}},\bar{Q}^{(\frac{1}{2},0)}_{r,\bar{r}},R^{(0,0)}_{r,\bar{r}}\}$:
$[M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{s}]=\left(\frac{1}{2}r-\frac{1}{2}s\right)Q^{(0,\frac{1}{2})}_{r+s,\bar{r}},\quad[M_{r,\bar{r}},\bar{Q}^{(0,\frac{1}{2})}_{\bar{s}}]=\left(\frac{1}{2}\bar{r}-\frac{1}{2}\bar{s}\right)\bar{Q}^{(\frac{1}{2},0)}_{r,\bar{r}+\bar{s}}$, (2.9ag)
$\{Q^{(\frac{1}{2},0)}_{r},Q^{(\frac{1}{2},0)}_{s}\}=L_{r+s},\quad\{Q^{(\frac{1}{2},0)}_{r},Q^{(0,\frac{1}{2})}_{s,\bar{s}}\}=M_{r+s,\bar{s}},$
$\{\bar{Q}^{(0,\frac{1}{2})}_{\bar{r}},\bar{Q}^{(0,\frac{1}{2})}_{\bar{s}}\}=\bar{L}_{\bar{r}+\bar{s}},\quad\{\bar{Q}^{(0,\frac{1}{2})}_{\bar{r}},\bar{Q}^{(\frac{1}{2},0)}_{s,\bar{s}}\}=M_{s,\bar{r}+\bar{s}},$
$\{Q^{(\frac{1}{2},0)}_{r},\bar{Q}^{(\frac{1}{2},0)}_{s,\bar{s}}\}=\left(\frac{1}{2}r-\frac{1}{2}s\right)R^{(0,0)}_{r+s,\bar{s}},\quad\{\bar{Q}^{(0,\frac{1}{2})}_{\bar{r}},Q^{(0,\frac{1}{2})}_{s,\bar{s}}\}=-\left(\frac{1}{2}\bar{r}-\frac{1}{2}\bar{s}\right)R^{(0,0)}_{s,\bar{r}+\bar{s}}$,
$[R^{(0,0)}_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{s}]=-\bar{Q}^{(\frac{1}{2},0)}_{r+s,\bar{s}},\quad[R^{(0,0)}_{r,\bar{r}},\bar{Q}^{(0,\frac{1}{2})}_{\bar{s}}]=Q^{(0,\frac{1}{2})}_{s,\bar{r}+\bar{s}}$,
Type II-II $\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{r},Q^{(1,\frac{1}{2})}_{r,\bar{r}},\bar{Q}^{(0,\frac{1}{2})}_{\bar{r}},\bar{Q}^{(\frac{1}{2},1)}_{r,\bar{r}},R^{(1,1)}_{r,\bar{r}}\}$:
$[M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{s}]=Q^{(1,\frac{1}{2})}_{r+s,\bar{r}},\quad[M_{r,\bar{r}},\bar{Q}^{(0,\frac{1}{2})}_{\bar{s}}]=\bar{Q}^{(\frac{1}{2},1)}_{r,\bar{r}+\bar{s}}$, (2.9ah)
$\{Q^{(\frac{1}{2},0)}_{r},Q^{(\frac{1}{2},0)}_{s}\}=L_{r+s},\quad\{Q^{(\frac{1}{2},0)}_{r},Q^{(1,\frac{1}{2})}_{s,\bar{s}}\}=-\left(r-\frac{1}{2}s\right)M_{r+s,\bar{s}},$
$\{\bar{Q}^{(0,\frac{1}{2})}_{\bar{r}},\bar{Q}^{(0,\frac{1}{2})}_{\bar{s}}\}=\bar{L}_{\bar{r}+\bar{s}},\quad\{\bar{Q}^{(0,\frac{1}{2})}_{\bar{r}},\bar{Q}^{(\frac{1}{2},1)}_{s,\bar{s}}\}=-\left(\bar{r}-\frac{1}{2}\bar{s}\right)M_{s,\bar{r}+\bar{s}},$
$\{Q^{(\frac{1}{2},0)}_{r},\bar{Q}^{(\frac{1}{2},1)}_{s,\bar{s}}\}=R^{(1,1)}_{r+s,\bar{s}},\quad\{\bar{Q}^{(0,\frac{1}{2})}_{\bar{r}},Q^{(1,\frac{1}{2})}_{s,\bar{s}}\}=-R^{(1,1)}_{s,\bar{r}+\bar{s}}$,
$[R^{(1,1)}_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{s}]=\left(\frac{1}{2}r-s\right)\bar{Q}^{(\frac{1}{2},1)}_{r+s,\bar{s}},\quad[R^{(1,1)}_{r,\bar{r}},\bar{Q}^{(0,\frac{1}{2})}_{\bar{s}}]=-\left(\frac{1}{2}\bar{r}-\bar{s}\right)Q^{(0,\frac{1}{2})}_{s,\bar{r}+\bar{s}}$.
Type I-II $\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{r},Q^{(0,\frac{1}{2})}_{r,\bar{r}},\bar{Q}^{(0,\frac{1}{2})}_{\bar{r}},\bar{Q}^{(\frac{1}{2},1)}_{r,\bar{r}},R^{(0,1)}_{r,\bar{r}}\}$:
$[M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{s}]=\left(\frac{1}{2}r-\frac{1}{2}s\right)Q^{(0,\frac{1}{2})}_{r+s,\bar{r}},\quad[M_{r,\bar{r}},\bar{Q}^{(0,\frac{1}{2})}_{\bar{s}}]=\bar{Q}^{(\frac{1}{2},1)}_{r,\bar{r}+\bar{s}}$, (2.9ai)
$\{Q^{(\frac{1}{2},0)}_{r},Q^{(\frac{1}{2},0)}_{s}\}=L_{r+s},\quad\{Q^{(\frac{1}{2},0)}_{r},Q^{(0,\frac{1}{2})}_{s,\bar{s}}\}=M_{r+s,\bar{s}},$
$\{\bar{Q}^{(0,\frac{1}{2})}_{\bar{r}},\bar{Q}^{(0,\frac{1}{2})}_{\bar{s}}\}=\bar{L}_{\bar{r}+\bar{s}},\quad\{\bar{Q}^{(0,\frac{1}{2})}_{\bar{r}},\bar{Q}^{(\frac{1}{2},1)}_{s,\bar{s}}\}=-\left(\bar{r}-\frac{1}{2}\bar{s}\right)M_{s,\bar{r}+\bar{s}},$
$\{Q^{(\frac{1}{2},0)}_{r},\bar{Q}^{(\frac{1}{2},1)}_{s,\bar{s}}\}=\left(\frac{1}{2}r-\frac{1}{2}s\right)R^{(0,1)}_{r+s,\bar{s}},\quad\{\bar{Q}^{(0,\frac{1}{2})}_{\bar{r}},Q^{(0,\frac{1}{2})}_{s,\bar{s}}\}=R^{(0,1)}_{s,\bar{r}+\bar{s}}$,
$[R^{(0,1)}_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{s}]=-\bar{Q}^{(\frac{1}{2},1)}_{r+s,\bar{s}},\quad[R^{(0,1)}_{r,\bar{r}},\bar{Q}^{(0,\frac{1}{2})}_{\bar{s}}]=\left(\frac{1}{2}\bar{r}-\bar{s}\right)Q^{(0,\frac{1}{2})}_{s,\bar{r}+\bar{s}}$,
as well as Type II-I, which is the conjugate of Type I-II in (2.9ai). Owing to this conjugation relation, we will omit the discussion of Type II-I, as it follows analogously from the Type I-II case.
For each case, we introduce distinct $R$-symmetry charges. In particular, $R^{(0,1)}$, $R^{(1,0)}$, and $R^{(1,1)}$ carry nonvanishing spin indices. This implies that, unlike the conventional $R$-symmetries which commute with spacetime symmetries, these $R$-symmetries transform covariantly under Lorentz rotations, distinguishing them from the usual extended super Poincaré algebras.
Finally, let us consider the case where two sets of $Q$s anticommute to $L$. To avoid introducing nonphysical $L_{n,\bar{n}}$ generators, the anti-holomorphic spins of these $Q$s must also vanish. Additionally, the spin-statistics requirement restricts us to two possibilities:
[…]
The first case reduces to the extended super BMS4 algebras of Type I (2.9ac), Type II (2.9ad), or their combinations, via appropriate rotations of $Q^{(\frac{1}{2},0)}_{r}$ and $\tilde{Q}^{(\frac{1}{2},0)}_{s}$. In the second case, an additional bosonic generator arising from $\{\tilde{Q}^{(\frac{3}{2},0)},\tilde{Q}^{(\frac{3}{2},0)}\}$ would be required to satisfy the Jacobi identity of $\{Q^{(\frac{1}{2},0)}_{r},\tilde{Q}^{(\frac{3}{2},0)},\tilde{Q}^{(\frac{3}{2},0)}\}$. However, this generator lacks a well-defined physical motivation. Hence, no unextended super BMS4 algebra exists with two sets of $Q$s anticommuting to $L$. The same argument applies to the anti-holomorphic sector.
To conclude, there are six unextended magnetic super BMS4 algebras: chiral Type I and Type II, as well as nonchiral Type I-I, Type I-II, Type II-I, and Type II-II. In these algebras, the finite-dimensional subalgebra of Type I contains only spin-$\frac{1}{2}$ supercharges. However, it violates the ansatz that the anticommutators of supercharges yield translations. Therefore, none of the magnetic superalgebras are connected to the super Poincaré algebra.
3 Free Field Realization
In this section, we present free field realizations of the various super BMS4 algebras.
3.1 Symmetry generators in terms of field modes
It is standard in QFT, particularly in 2D free CFTs and in BMS3 and BMS4 free theories, to express symmetry generators in terms of field modes [60, 61, 62, 63, 64, 65, 66]. The conventional approach to obtain the symmetry generators of a given free theory involves first computing the energy-momentum tensor in terms of the fields, then performing a Laurent or Fourier expansion of the fields according to their equations of motion (EOMs), and finally substituting this expansion to read off the generators. Here, since the symmetry algebra is already known, we adopt an inverse approach: we construct the symmetry generators using a specific ansatz for the field modes, and then build the corresponding field theory consistent with these mode expansions.
The expansion of fields is analogous to the expansion (2.2) of conserved currents. For example, a time independent field is expanded as
$\phi(t,z,\bar{z})=\sum_{r,\bar{r}}\phi^{(h,\bar{h})}_{r,\bar{r}}z^{-r-h}\bar{z}^{-\bar{r}-\bar{h}}.$ (2.9a)
Consequently, the modes of free fields $\phi^{(h,\bar{h})}_{r,\bar{r}}$ also fall into the charge representations discussed in Section 2.1. For simplicity, we show only the holomorphic part:
$L_{m},\phi^{(h)}_{r}]=((h-1)m-r)\phi^{(h)}_{m+r}$, (2.9b)
Similarly, we label the representations by l=h1l=h-1. Assuming that bilinears of modes $\phi^{i}$ form a set of charges $Q$ in the $(l_{Q})$ representation,
$Q^{(l_{Q})}_{m}=\sum_{r}c^{l_{Q}}_{l_{i},l_{j}}(r,n)\phi^{i}_{r}\phi^{j}_{m-r},$ (2.9c)
the consistency condition from the commutator $L_{m},Q]$ imposes a constraint on the function $c$:
$((l_{i}+1)m-r)c(r-m,t)+(l_{j}m-t+r)c(r,t)-(l_{Q}m-t)c(r,t+m)=0.$ (2.9d)
Note that this equation for $c$ differs from the equation for $f$ in (2.8). For constructing free theories, it suffices to consider the following solutions:
$c_{l_{i},l_{j}}^{l_{Q}}(r,n)=1,$ $l_{Q}=l_{i}+l_{j}+1,$ (2.9ea)
$c_{l_{i},l_{j}}^{l_{Q}}(r,n)=l_{Q}r-(l_{i}+1)n,$ $l_{Q}=l_{i}+l_{j}+2,$ (2.9eb)
Other solutions can be found in Appendix A.
In the free theory, another crucial point is the canonical quantization. Suppose $\varphi$ and $\phi$ are two sets of canonically conjugate modes. Their commutation relation is expected to take the form
$\varphi_{r},\phi_{s}\}=g^{l_{I}}_{l_{\varphi},l_{\phi}}(r,s)\delta_{r+s}I$ (2.9f)
where $\delta$ denotes the Kronecker delta and $I$ is the identity operator. In particular, the identity generator $I$ is regarded as a charge in the representation with (anti-)holomorphic labels $l_{I}=\bar{l}_{I}=0$ and (anti-)holomorphic indices $m_{I}=\bar{m}_{I}=0$. Indeed, if we formally extend $I$ to $I_{m}$, we have
$L_{m},I_{n=0}]=(l_{I}*m-1*n)I_{m}=(0*m-1*0)I_{m}=0,$ (2.9g)
showing that additional generators $I_{m\neq 0}$ do not enter the theory as symmetry generators. Therefore, $\varphi$, $\phi$, and $I$ can all be viewed as charges in the theory, and equation (2.9f) implies that $g(r,s)\delta_{r+s}$ satisfies the constraint (2.8). The discrete solution relevant for this situation is
$f^{l_{I}}_{l_{\varphi},l_{\phi}}(r,s)=\delta_{r+s},\quad l_{\varphi}+l_{\phi}=-1,\quad l_{I}=0.$ (2.9h)
As an example, consider the BMS4 algebra. Suppose the theory contains bosonic modes $\varphi^{(-\frac{3}{4},-\frac{3}{4})}$ and $\phi^{(-\frac{1}{4},-\frac{1}{4})}$, which are canonically conjugate to each other. Their commutation relation reads
$\varphi^{(-\frac{3}{4},-\frac{3}{4})}_{r,\bar{r}},\phi^{(-\frac{1}{4},-\frac{1}{4})}_{s,\bar{s}}]=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}},$ (2.9i)
where the identity operator is implicit. Using these modes, we can construct the BMS4 algebra as
$L_{m}=-4\pi^{2}i\sum_{s,\bar{s}}\left(s-\frac{1}{4}m\right)\varphi_{s,\bar{s}}\phi_{m-s,-\bar{s}},$ (2.9j)
$\bar{L}_{\bar{m}}=-4\pi^{2}i\sum_{s,\bar{s}}\left(\bar{s}-\frac{1}{4}\bar{m}\right)\varphi_{s,\bar{s}}\phi_{-s,\bar{m}-\bar{s}},$
$M_{r,\bar{r}}=2\pi^{2}i\sum_{s,\bar{s}}\phi_{s,\bar{s}}\phi_{r-s,\bar{r}-\bar{s}}.$
In this construction, rule (2.9ea) is satisfied by the anti-holomorphic index of $L$, the holomorphic index of $\bar{L}$, and both indices of $M$, while rule (2.9eb) is satisfied by the holomorphic index of $L$ and the anti-holomorphic index of $\bar{L}$. In defining $L$ and $\bar{L}$, we consistently order the modes such that $\phi$ is always to the right of $\varphi$. This reproduces the BMS4 algebra (1.1) without introducing central terms. The ordering convention will be implicitly applied throughout the rest of this section. Furthermore, using the commutation relation (2.9i), one can verify that the transformation of the modes under the BMS4 algebra:
$[L_{m},\varphi^{(-\frac{3}{4},-\frac{3}{4})}_{s,\bar{s}}]=\left(-\frac{3}{4}m-s\right)\varphi^{(-\frac{3}{4},-\frac{3}{4})}_{m+s,\bar{s}}$, $[L_{m},\phi^{(-\frac{1}{4},-\frac{1}{4})}_{s,\bar{s}}]=\left(-\frac{1}{4}m-s\right)\phi^{(-\frac{1}{4},-\frac{1}{4})}_{m+s,\bar{s}}$, (2.9k)
$[\bar{L}_{\bar{m}},\varphi^{(-\frac{3}{4},-\frac{3}{4})}_{s,\bar{s}}]=\left(-\frac{3}{4}\bar{m}-\bar{s}\right)\varphi^{(-\frac{3}{4},-\frac{3}{4})}_{s,\bar{m}+\bar{s}}$, $[\bar{L}_{\bar{m}},\phi^{(-\frac{1}{4},-\frac{1}{4})}_{s,\bar{s}}]=\left(-\frac{1}{4}\bar{m}-\bar{s}\right)\phi^{(-\frac{1}{4},-\frac{1}{4})}_{s,\bar{m}+\bar{s}}$,
$[M_{r,\bar{r}},\varphi^{(-\frac{3}{4},-\frac{3}{4})}_{s,\bar{s}}]=\phi^{(-\frac{1}{4},-\frac{1}{4})}_{r+s,\bar{r}+\bar{s}}$, $[M_{r,\bar{r}},\phi^{(-\frac{1}{4},-\frac{1}{4})}_{s,\bar{s}}]=0.$
These commutation relations satisfy (2.9a) and (2.9b) up to overall factors respectively. In this example, negative spin indices are allowed because the existence of finite-dimensional subsectors is not essential for field modes.
Consider the BMS4 theory on $\mathbb{R}\times\mathbb{T}^{2}$, where the coordinates $(t,\sigma,\bar{\sigma})$ are reparametrized as $(t,z,\bar{z})$ with $z=e^{i\sigma}$ and $\bar{z}=e^{i\bar{\sigma}}$. The BMS4 algebra is generated by the following vector fields:
$\xi^{(L)}_{n}=\left(-\frac{n+1}{2}tz^{n},-z^{(n+1)},0\right),$ (2.9l)
$\xi^{(\bar{L})}_{\bar{n}}=\left(-\frac{\bar{n}+1}{2}t\bar{z}^{n},0,-\bar{z}^{(\bar{n}+1)}\right),$
$\xi^{(M)}_{r,\bar{r}}=\left(-z^{r}\bar{z}^{\bar{r}},0,0\right).$
For a BMS4 field $\phi$, the action of the symmetry transformation $G_{\xi}$ generated by $\xi^{\mu}$ is given by [67]
$[L_{n},\phi]=-(\xi^{(L)}_{n})^{\mu}\partial_{\mu}\phi-2h\partial_{t}(\xi^{(L)}_{n})^{t}\phi+\frac{1}{2}\partial_{z}(\xi^{(L)}_{n})^{t}\phi_{z}+\frac{1}{2}\partial_{\bar{z}}(\xi^{(L)}_{n})^{t}\phi_{\bar{z}}+\cdots,$ (2.9m)
$[\bar{L}_{\bar{n}},\phi]=-(\xi^{(\bar{L})}_{\bar{n}})^{\mu}\partial_{\mu}\phi-2\bar{h}\partial_{t}(\xi^{(\bar{L})}_{\bar{n}})^{t}\phi+\frac{1}{2}\partial_{z}(\xi^{(\bar{L})}_{\bar{n}})^{t}\phi_{z}+\frac{1}{2}\partial_{\bar{z}}(\xi^{(\bar{L})}_{\bar{n}})^{t}\phi_{\bar{z}}+\cdots,$
$[M_{r,\bar{r}},\phi]=-(\xi^{(M)}_{r,\bar{r}})^{\mu}\partial_{\mu}\phi+\frac{1}{2}\partial_{z}(\xi^{(M)}_{r,\bar{r}})^{t}\phi_{z}+\frac{1}{2}\partial_{\bar{z}}(\xi^{(M)}_{r,\bar{r}})^{t}\phi_{\bar{z}}+\cdots.$
Here, the fields $\phi_{z}$ and $\phi_{\bar{z}}$ belong to the same Carrollian multiplet as the field $\phi$ [68], while the ellipses denote contributions from higher-order multiplet fields. The structure of this relation is illustrated in Figure 1.
Figure 1: Structure of Carrollian multiplet of $\phi$. The ellipses denote higher-order multiplet fields.
To meet equations (2.9m), the example discussed here can be identified as the free electric BMS4 scalar theory:
$\mathcal{L}=\frac{1}{2}\partial_{0}\phi\partial_{0}\phi.$ (2.9n)
The EOM of the field $\phi$ is $\partial_{0}^{2}\phi=0$, and $\phi$ can be expanded as
$\phi(t,z,\bar{z})=\sum_{r,\bar{r}}\varphi_{r,\bar{r}}z^{-r-\frac{1}{4}}\bar{z}^{-\bar{r}-\frac{1}{4}}+t\phi_{r,\bar{r}}z^{-r-\frac{3}{4}}\bar{z}^{-\bar{r}-\frac{3}{4}}.$ (2.9o)
In this construction, the conformal dimensions of the field $\phi$ are $h=\bar{h}=\frac{1}{4}$, giving $\Delta=\frac{1}{2}$ and spin $j=0$. The commutation relation (2.9i) is consistent with the canonical commutation relation of the theory:
$\phi(t,z,\bar{z}),\partial_{0}\phi(t,z^{\prime},\bar{z}^{\prime})]=i\delta^{2}(z-z^{\prime}).$ (2.9p)
In the following subsections, we use the same method to construct supersymmetric BMS4 theories with the symmetries discussed in Section 2. For simplicity, we focus on scalar multiplets of different supersymmetries, with the scalar theory taken as (2.9n).
3.2 Electric super BMS4 algebra
In Section 2.2, we identified ten electric super BMS4 algebras (2.9s)–(2.9ab). However, in the free field mode construction, only the algebras (2.9s), (2.9t), and (2.9u) admit supersymmetry between spin-0 scalars and spin-$\frac{1}{2}$ fermions. Realizing the other algebras requires introducing higher-spin fermion fields. Since, in the bulk 4D unextended super Poincaré theory, the superpartner of a scalar is a $\frac{1}{2} spin$or, we focus here only on the free field realizations of (2.9s), (2.9t), and (2.9u).
Realization of electric algebra (2.9s)
To realize (2.9s), we introduce bosonic field modes
$\varphi^{(-\frac{3}{4},-\frac{3}{4})},\quad\phi^{(-\frac{1}{4},-\frac{1}{4})},\quad\text{with }[\varphi^{(-\frac{3}{4},-\frac{3}{4})}_{r,\bar{r}},\phi^{(-\frac{1}{4},-\frac{1}{4})}_{s,\bar{s}}]=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}},$ (2.9q)
and fermionic field modes
$\psi^{(-\frac{1}{4},-\frac{3}{4})},\quad\bar{\psi}^{(-\frac{3}{4},-\frac{1}{4})},\quad\text{with }\{\psi^{(-\frac{1}{4},-\frac{3}{4})}_{r,\bar{r}},\bar{\psi}^{(-\frac{3}{4},-\frac{1}{4})}_{s,\bar{s}}\}=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}}.$ (2.9r)
The super BMS4 algebra (2.9s) can then be realized as
$L_{m}=-4\pi^{2}i\sum_{s,\bar{s}}\left(s-\frac{1}{4}m\right)\varphi_{s,\bar{s}}\phi_{m-s,-\bar{s}}-\left(s-\frac{3}{4}m\right)\psi_{s,\bar{s}}\bar{\psi}_{m-s,-\bar{s}},$ (2.9s)
$\bar{L}_{\bar{m}}=-4\pi^{2}i\sum_{s,\bar{s}}\left(\bar{s}-\frac{1}{4}\bar{m}\right)\varphi_{s,\bar{s}}\phi_{-s,\bar{m}-\bar{s}}-\left(\bar{s}-\frac{1}{4}\bar{m}\right)\psi_{s,\bar{s}}\bar{\psi}_{-s,\bar{m}-\bar{s}},$
$M_{r,\bar{r}}=2\pi^{2}i\sum_{s,\bar{s}}\phi_{s,\bar{s}}\phi_{r-s,\bar{r}-\bar{s}},$
together with the supercharges
$Q^{(\frac{1}{2},0)}_{r,\bar{r}}=2\sqrt{2}\pi^{2}\sum_{s,\bar{s}}\phi_{s,\bar{s}}\psi_{r-s,\bar{r}-\bar{s}},\quad Q^{(0,\frac{1}{2})}_{r,\bar{r}}=2\sqrt{2}\pi^{2}\sum_{s,\bar{s}}\phi_{s,\bar{s}}\bar{\psi}_{r-s,\bar{r}-\bar{s}}.$ (2.9t)
It is straightforward for the interested reader to check that these constructions satisfy the commutation relations (2.9s) using (2.9q) and (2.9r). The transformations of the field modes under the super BMS4 algebra are summarized in Table 1. For simplicity, the actions of $L$ and $\bar{L}$ are omitted in the table, as the (anti-)holomorphic indices already encode their transformations.
Table 1: Field modes transformation under super BMS4 algebra (2.9s).
$\varphi_{s,\bar{s}}$ $\phi_{s,\bar{s}}$ $\psi_{s,\bar{s}}$ $\bar{\psi}_{s,\bar{s}}$
$M_{r,\bar{r}}$ $\phi_{r+s,\bar{r}+\bar{s}}$ 0 0 0
$Q^{(\frac{1}{2},0)}_{r,\bar{r}}$ $\frac{i}{\sqrt{2}}\psi_{r+s,\bar{r}+\bar{s}}$ 0 0 $\frac{i}{\sqrt{2}}\phi_{r+s,\bar{r}+\bar{s}}$
$Q^{(0,\frac{1}{2})}_{r,\bar{r}}$ $\frac{i}{\sqrt{2}}\bar{\psi}_{r+s,\bar{r}+\bar{s}}$ 0 $\frac{i}{\sqrt{2}}\phi_{r+s,\bar{r}+\bar{s}}$ 0
The corresponding field theory can be written as
$\mathcal{L}=\frac{1}{2}\partial_{0}\phi\partial_{0}\phi-\psi\partial_{0}\bar{\psi}-\bar{\psi}\partial_{0}\psi,$ (2.9y)
where the bosonic part is exactly (2.9n), and $\psi$ and $\bar{\psi}$ are fermionic fields with spin $\pm\frac{1}{2}$ respectively. In terms of Carrollian theories, $\phi$ is also referred as electric scalar, and $\psi$ are referred as electric spinor. The conformal dimensions are $\Delta_{\phi}=\frac{1}{2}$ and $\Delta_{\psi}=\Delta_{\bar{\psi}}=1$. The fermionic fields satisfy the equations of motion
$\partial_{0}\psi=0,\quad\partial_{0}\bar{\psi}=0,$ (2.9z)
with mode expansions
$\psi(t,z,\bar{z})=\sum_{r,\bar{r}}\psi_{r,\bar{r}}z^{-r-\frac{3}{4}}\bar{z}^{-\bar{r}-\frac{1}{4}},\quad\bar{\psi}(t,z,\bar{z})=\sum_{r,\bar{r}}\bar{\psi}_{r,\bar{r}}z^{-r-\frac{1}{4}}\bar{z}^{-\bar{r}-\frac{3}{4}}.$ (2.9aa)
The canonical anticommutation relation
$\{\phi(t,z,\bar{z}),\bar{\psi}(t,z^{\prime},\bar{z}^{\prime})\}=i\delta^{2}(z-z^{\prime})$ (2.9ab)
is consistent with the mode commutator (2.9r). In this theory, the bosonic field $\phi$ has one degree of freedom, while $\psi$ and $\bar{\psi}$ together account for one fermionic degree of freedom, as they are canonical conjugates. Hence, the numbers of bosonic and fermionic degrees of freedom are equal. Similarly, in the other constructed theories, the total bosonic and fermionic degrees of freedom match, even if the number of individual fields differs.
Realization of electric algebra (2.9t)
The same bosonic modes (2.9q) and fermionic modes (2.9r) can be used to realize (2.9t). The construction of the BMS4 algebra remains the same as in (2.9s), but the supercharges are now defined differently:
$Q^{(\frac{1}{2},1)}_{r,\bar{r}}=2\sqrt{2}\pi^{2}\sum_{s,\bar{s}}\left(\bar{s}-\frac{3}{4}\bar{r}\right)\phi_{s,\bar{s}}\psi_{r-s,\bar{r}-\bar{s}},\quad Q^{(0,\frac{1}{2})}_{r,\bar{r}}=2\sqrt{2}\pi^{2}\sum_{s,\bar{s}}\phi_{s,\bar{s}}\bar{\psi}_{r-s,\bar{r}-\bar{s}}.$ (2.9ac)
The transformation of the field modes under the bosonic symmetries is the same as in Table 1, while their transformation under the supercharges is summarized in Table 2.
Table 2: Field modes transformation under supercharges of (2.9t).
$\varphi_{s,\bar{s}}$ $\phi_{s,\bar{s}}$ $\psi_{s,\bar{s}}$ $\bar{\psi}_{s,\bar{s}}$
$Q^{(\frac{1}{2},1)}_{r,\bar{r}}$ $\frac{i}{\sqrt{2}}\left(-\frac{3}{4}\bar{r}-\bar{s}\right)\psi_{r+s,\bar{r}+\bar{s}}$ 0 0 $\frac{i}{\sqrt{2}}\left(-\frac{1}{4}\bar{r}-\bar{s}\right)\phi_{r+s,\bar{r}+\bar{s}}$
$Q^{(0,\frac{1}{2})}_{r,\bar{r}}$ $\frac{i}{\sqrt{2}}\bar{\psi}_{r+s,\bar{r}+\bar{s}}$ 0 $\frac{i}{\sqrt{2}}\phi_{r+s,\bar{r}+\bar{s}}$ 0
Realization of electric algebra (2.9u)
The same bosonic and fermionic modes (2.9q) and (2.9r) can also be employed to realize (2.9u), with a distinct construction for the supercharges:
$Q^{(\frac{1}{2},0)}_{r,\bar{r}}=2\sqrt{2}\pi^{2}\sum_{s,\bar{s}}\phi_{s,\bar{s}}\psi_{r-s,\bar{r}-\bar{s}},\quad Q^{(0,\frac{3}{2})}_{r,\bar{r}}=2\sqrt{2}\pi^{2}\sum_{s,\bar{s}}\left(\frac{3}{2}\bar{s}-\frac{3}{4}\bar{r}\right)\phi_{s,\bar{s}}\bar{\psi}_{r-s,\bar{r}-\bar{s}}.$ (2.9ag)
The transformations of the field modes under these supercharges are summarized in Table 3.
Table 3: Field modes transformation under supercharges of (2.9u).
$\varphi_{s,\bar{s}}$ $\phi_{s,\bar{s}}$ $\psi_{s,\bar{s}}$ $\bar{\psi}_{s,\bar{s}}$
$Q^{(\frac{1}{2},0)}_{r,\bar{r}}$ $\frac{i}{\sqrt{2}}\psi_{r+s,\bar{r}+\bar{s}}$ 0 0 $\frac{i}{\sqrt{2}}\phi_{r+s,\bar{r}+\bar{s}}$
$Q^{(0,\frac{3}{2})}_{r,\bar{r}}$ $\frac{i}{\sqrt{2}}\left(-\frac{3}{4}\bar{r}-\frac{3}{2}\bar{s}\right)\bar{\psi}_{r+s,\bar{r}+\bar{s}}$ 0 $\frac{i}{\sqrt{2}}\left(-\frac{3}{4}\bar{r}-\frac{3}{2}\bar{s}\right)\phi_{r+s,\bar{r}+\bar{s}}$ 0
Interestingly, the theory (2.9y) is sufficiently simple that it accommodates a large number of supercharges. Explicitly counting the conjugation, the field modes in (2.9y) can generate the following supercharges:
$Q^{(\frac{1}{2},0)},\quad Q^{(0,\frac{1}{2})},\quad Q^{(\frac{1}{2},1)},\quad Q^{(1,\frac{1}{2})},\quad Q^{(\frac{3}{2},0)},\quad Q^{(0,\frac{3}{2})},$ (2.9ak)
where all except $Q^{(1,\frac{1}{2})}$ and $Q^{(\frac{3}{2},1)}$ were constructed previously. However, considering all supercharges simultaneously quickly becomes cumbersome. For instance, one finds that
$\{Q^{(\frac{1}{2},1)}_{r,\bar{r}},Q^{(0,\frac{3}{2})}_{s,\bar{s}}\}=-i(4\pi^{2})\sum_{t,\bar{t}}-\frac{3}{16}(3\bar{r}^{2}+2\bar{r}(\bar{s}-4\bar{t})+2(\bar{s}-2\bar{t})^{2})\phi_{t,\bar{t}}\phi_{r+s-t,\bar{r}+\bar{s}-\bar{t}}.$ (2.9al)
This shows that the anticommutator introduces bosonic generators beyond $M$. Similar structures appear in other anticommutators as well. While these additional generators are indeed symmetries of the specific theory (2.9y), they do not correspond to symmetries of generic BMS4 theories.
3.3 Magnetic super BMS4 algebra
We now turn to the realization of the magnetic super BMS4 algebras. The Type I algebra (2.9ac) can be constructed using the same modes as in (2.9q) and (2.9r). However, the altered transformation rules under the symmetry imply that these modes correspond to a different physical theory. In contrast, realizing the Type II algebra (2.9ad) requires introducing a distinct set of modes.
Realization of magnetic Type I algebra (2.9ac)
For convenience, we repeat the modes (2.9q) and (2.9r) here, with a renaming of one set of fermionic modes:
$\varphi^{(-\frac{3}{4},-\frac{3}{4})},\quad\phi^{(-\frac{1}{4},-\frac{1}{4})},\quad[\varphi^{(-\frac{3}{4},-\frac{3}{4})}_{r,\bar{r}},\phi^{(-\frac{1}{4},-\frac{1}{4})}_{s,\bar{s}}]=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}},$ (2.9am)
$\psi^{(-\frac{1}{4},-\frac{3}{4})},\quad\chi^{(-\frac{3}{4},-\frac{1}{4})},\quad\{\psi^{(-\frac{1}{4},-\frac{3}{4})}_{r,\bar{r}},\chi^{(-\frac{3}{4},-\frac{1}{4})}_{s,\bar{s}}\}=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}}.$
The Type I algebra (2.9ac) is realized as
$L_{m}=-4\pi^{2}i\sum_{s,\bar{s}}\left(s-\frac{1}{4}m\right)\varphi_{s,\bar{s}}\phi_{m-s,-\bar{s}}-\left(s-\frac{3}{4}m\right)\psi_{s,\bar{s}}\chi_{m-s,-\bar{s}},$ (2.9an)
$\bar{L}_{\bar{m}}=-4\pi^{2}i\sum_{s,\bar{s}}\left(\bar{s}-\frac{1}{4}\bar{m}\right)\varphi_{s,\bar{s}}\phi_{-s,\bar{m}-\bar{s}}-\left(\bar{s}-\frac{1}{4}\bar{m}\right)\psi_{s,\bar{s}}\chi_{-s,\bar{m}-\bar{s}},$
$M_{r,\bar{r}}=2\pi^{2}i\sum_{s,\bar{s}}\phi_{s,\bar{s}}\phi_{r-s,\bar{r}-\bar{s}}+2\left(\frac{1}{2}s-\frac{1}{4}r\right)\chi_{s,\bar{s}}\chi_{r-s,\bar{r}-\bar{s}},$
and
$Q^{(\frac{1}{2},0)}_{r}=-2\sqrt{2}\pi^{2}i\sum_{s,\bar{s}}2\left(\frac{1}{2}s-\frac{1}{4}r\right)\varphi_{s,\bar{s}}\chi_{r-s,-\bar{s}}+\phi_{s,\bar{s}}\psi_{r-s,-\bar{s}},$ (2.9ao)
$Q^{(0,\frac{1}{2})}_{r,\bar{r}}=2\sqrt{2}\pi^{2}i\sum_{s,\bar{s}}\phi_{s,\bar{s}}\chi_{r-s,\bar{r}-\bar{s}}.$
Notice that the construction of $M$ in (2.9an) differs from that in (2.9s). Under the algebra (2.9ac), the modes transform as summarized in Table 4.
Table 4: Field modes transformation under Type I BMS4 algebra (2.9ac).
$\varphi_{s,\bar{s}}$ $\phi_{s,\bar{s}}$ $\psi_{s,\bar{s}}$ $\chi_{s,\bar{s}}$
$M_{r,\bar{r}}$ $\phi_{r+s,\bar{r}+\bar{s}}$ 0 $2\left(\frac{1}{4}r+\frac{1}{2}s\right)\chi_{r+s,\bar{r}+\bar{s}}$ 0
$Q^{(\frac{1}{2},0)}_{r}$ $\frac{1}{\sqrt{2}}\psi_{r+s,\bar{s}}$ $\sqrt{2}\left(-\frac{1}{4}r-\frac{1}{2}s\right)\chi_{r+s,\bar{s}}$ $\sqrt{2}\left(-\frac{1}{4}r-\frac{1}{2}s\right)\varphi_{r+s,\bar{s}}$ $\frac{1}{\sqrt{2}}\phi_{r+s,\bar{s}}$
$Q^{(0,\frac{1}{2})}_{r,\bar{r}}$ $\frac{1}{\sqrt{2}}\chi_{r+s,\bar{r}+\bar{s}}$ 0 $\frac{1}{\sqrt{2}}\phi_{r+s,\bar{r}+\bar{s}}$ 0
The corresponding field theory for this construction is
$\mathcal{L}=\frac{1}{2}\partial_{0}\phi\partial_{0}\phi-\psi_{+}\partial_{0}\psi_{-}-\psi_{-}\partial_{0}\psi_{+}+\psi_{-}\partial_{z}\psi_{-}.$ (2.9at)
In this theory, there is a single scalar field $\phi$ with conformal dimension $\Delta_{\phi}=\frac{1}{2}$, and two fermionic fields $\psi_{\pm}$ with spin $\pm\frac{1}{2}$ and conformal dimensions $\Delta_{\psi_{\pm}}=1$. The equations of motion derived from the Lagrangian read
$\partial_{0}^{2}\phi=0,\quad\partial_{0}\psi_{+}=-\partial_{z}\psi_{-},\quad\partial_{0}\psi_{-}=0,$ (2.9au)
and the corresponding mode expansions consistent with these EOMs are
[…]
Realization of magnetic Type II algebra (2.9ad)
The construction of the magnetic Type II algebra employs slightly different modes. The bosonic modes remain the same as in the Type I case, while the fermionic modes carry different (anti-)holomorphic spins:
$\varphi^{(-\frac{3}{4},-\frac{3}{4})},\quad\phi^{(-\frac{1}{4},-\frac{1}{4})},\quad[\varphi^{(-\frac{3}{4},-\frac{3}{4})}_{r,\bar{r}},\phi^{(-\frac{1}{4},-\frac{1}{4})}_{s,\bar{s}}]=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}},$ (2.9ax)
$\psi^{(-\frac{5}{4},-\frac{3}{4})},\quad\chi^{(\frac{1}{4},-\frac{1}{4})},\quad\{\psi^{(-\frac{5}{4},-\frac{3}{4})}_{r,\bar{r}},\chi^{(\frac{1}{4},-\frac{1}{4})}_{s,\bar{s}}\}=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}}.$
The Type II algebra (2.9ad) is realized through the following mode constructions:
$L_{m}=-4\pi^{2}i\sum_{s,\bar{s}}\left(s-\frac{1}{4}m\right)\varphi_{s,\bar{s}}\phi_{m-s,-\bar{s}}-\left(s+\frac{1}{4}m\right)\psi_{s,\bar{s}}\chi_{m-s,-\bar{s}},$ (2.9ay)
$\bar{L}_{\bar{m}}=-4\pi^{2}i\sum_{s,\bar{s}}\left(\bar{s}-\frac{1}{4}\bar{m}\right)\varphi_{s,\bar{s}}\phi_{-s,\bar{m}-\bar{s}}-\left(\bar{s}-\frac{1}{4}\bar{m}\right)\psi_{s,\bar{s}}\chi_{-s,\bar{m}-\bar{s}},$
$M_{r,\bar{r}}=2\pi^{2}i\sum_{s,\bar{s}}\phi_{s,\bar{s}}\phi_{r-s,\bar{r}-\bar{s}},$
$Q^{(\frac{1}{2},0)}_{r}=-4\pi^{2}i\sum_{s,\bar{s}}\varphi_{s,\bar{s}}\chi_{r-s,-\bar{s}}-\left(\frac{1}{2}s-\frac{3}{4}r\right)\phi_{s,\bar{s}}\psi_{r-s,-\bar{s}},$ (2.9az)
$Q^{(1,\frac{1}{2})}_{r,\bar{r}}=-4\pi^{2}i\sum_{s,\bar{s}}\phi_{s,\bar{s}}\chi_{r-s,\bar{r}-\bar{s}}.$
The transformations of these modes under the Type II algebra are summarized in Table 5.
Table 5: Field modes transformation under Type II BMS4 algebra (2.9ad).
$\varphi_{s,\bar{s}}$ $\phi_{s,\bar{s}}$ $\psi_{s,\bar{s}}$ $\chi_{s,\bar{s}}$
$M_{r,\bar{r}}$ $\phi_{r+s,\bar{r}+\bar{s}}$ 0 0 0
$Q^{(\frac{1}{2},0)}_{r}$ $\left(-\frac{3}{4}r-\frac{1}{2}s\right)\psi_{r+s,\bar{s}}$ $\chi_{r+s,\bar{s}}$ $\varphi_{r+s,\bar{s}}$ $\left(\frac{1}{4}r-\frac{1}{2}s\right)\phi_{r+s,\bar{s}}$
$Q^{(0,\frac{1}{2})}_{r,\bar{r}}$ $\chi_{r+s,\bar{r}+\bar{s}}$ 0 $\phi_{r+s,\bar{r}+\bar{s}}$ 0
The field theory corresponding to this construction is given by
$\mathcal{L}=\frac{1}{2}\partial_{0}\phi\partial_{0}\phi-\psi_{+}\partial_{0}\psi_{-}-\psi_{-}\partial_{0}\psi_{+}.$ (2.9be)
This theory consists of a single scalar field $\phi$ and two fermionic fields $\psi_{\pm}$ with spins $\pm\frac{1}{2}$. Although the Lagrangian (2.9be) appears similar to the electric super BMS4 field theory (2.9y), it describes a different theory because the conformal dimensions of the fermionic fields differ: $\Delta_{\psi_{+}}=2$ and $\Delta_{\psi_{-}}=0$. The equations of motion read
$\partial_{0}^{2}\phi=0,\quad\partial_{0}\psi_{+}=\partial_{0}\psi_{-}=0,$ (2.9bf)
and the consistent mode expansions for these fields are
[…]
Realization of magnetic Type I-I algebra (2.9ag)
We now consider a theory that possesses both left- and right-handed super BMS4 symmetries. The mode content of this theory is specified as follows:
$\varphi^{(-\frac{3}{4},-\frac{3}{4})},\quad\phi^{(-\frac{1}{4},-\frac{1}{4})},\quad[\varphi^{(-\frac{3}{4},-\frac{3}{4})}_{r,\bar{r}},\phi^{(-\frac{1}{4},-\frac{1}{4})}_{s,\bar{s}}]=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}},$ (2.9bi)
$\psi^{(-\frac{1}{4},-\frac{3}{4})},\quad\chi^{(-\frac{3}{4},-\frac{1}{4})},\quad\{\psi^{(-\frac{1}{4},-\frac{3}{4})}_{r,\bar{r}},\chi^{(-\frac{3}{4},-\frac{1}{4})}_{s,\bar{s}}\}=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}},$
$\bar{\psi}^{(-\frac{3}{4},-\frac{1}{4})},\quad\bar{\chi}^{(-\frac{1}{4},-\frac{3}{4})},\quad\{\bar{\psi}^{(-\frac{3}{4},-\frac{1}{4})}_{r,\bar{r}},\bar{\chi}^{(-\frac{1}{4},-\frac{3}{4})}_{s,\bar{s}}\}=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}},$
$\bar{\varphi}^{(-\frac{1}{4},-\frac{1}{4})},\quad\bar{\phi}^{(-\frac{3}{4},-\frac{3}{4})},\quad[\bar{\varphi}^{(-\frac{1}{4},-\frac{1}{4})}_{r,\bar{r}},\bar{\phi}^{(-\frac{3}{4},-\frac{3}{4})}_{s,\bar{s}}]=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}}.$
The Type I-I algebra (2.9ag) features an additional bosonic generator $R$. The BMS4 part of the algebra can be realized as
$L_{m}=-4\pi^{2}i\sum_{s,\bar{s}}\left(s-\frac{1}{4}m\right)\varphi_{s,\bar{s}}\phi_{m-s,-\bar{s}}-\left(s-\frac{3}{4}m\right)\psi_{s,\bar{s}}\chi_{m-s,-\bar{s}}$ (2.9bj)
$\qquad\qquad\qquad\qquad-\left(s-\frac{1}{4}m\right)\bar{\psi}_{s,\bar{s}}\bar{\chi}_{m-s,-\bar{s}}+\left(s-\frac{3}{4}m\right)\bar{\varphi}_{s,\bar{s}}\bar{\phi}_{m-s,-\bar{s}},$
$\bar{L}_{\bar{m}}=-4\pi^{2}i\sum_{s,\bar{s}}\left(\bar{s}-\frac{1}{4}\bar{m}\right)\varphi_{s,\bar{s}}\phi_{-s,\bar{m}-\bar{s}}-\left(\bar{s}-\frac{1}{4}\bar{m}\right)\psi_{s,\bar{s}}\chi_{-s,\bar{m}-\bar{s}}$
$\qquad\qquad\qquad\qquad-\left(\bar{s}-\frac{3}{4}\bar{m}\right)\bar{\psi}_{s,\bar{s}}\bar{\chi}_{-s,\bar{m}-\bar{s}}+\left(\bar{s}-\frac{3}{4}\bar{m}\right)\bar{\varphi}_{s,\bar{s}}\bar{\phi}_{-s,\bar{m}-\bar{s}},$
$M_{r,\bar{r}}=2\pi^{2}i\sum_{s,\bar{s}}\phi_{s,\bar{s}}\phi_{r-s,\bar{r}-\bar{s}}+2\left(\frac{1}{2}\bar{s}-\frac{1}{4}\bar{m}\right)\chi_{s,\bar{s}}\chi_{r-s,\bar{m}-\bar{s}}$
$\qquad\qquad+2\left(\frac{1}{2}\bar{s}-\frac{1}{4}\bar{r}\right)\bar{\chi}_{s,\bar{s}}\bar{\chi}_{m-s,\bar{r}-\bar{s}}-2\left(\frac{1}{2}\bar{s}-\frac{1}{4}\bar{m}\right)\left(\frac{1}{2}\bar{s}-\frac{1}{4}\bar{r}\right)\bar{\phi}_{s,\bar{s}}\bar{\phi}_{m-s,\bar{r}-\bar{s}},$
the fermionic generators are constructed as
$Q^{(\frac{1}{2},0)}_{r}=-2\pi^{2}i\sum_{s,\bar{s}}2\sqrt{2}\left(\frac{1}{2}s-\frac{1}{4}r\right)\varphi_{s,\bar{s}}\chi_{r-s,-\bar{s}}+\sqrt{2}~\phi_{s,\bar{s}}\psi_{r-s,-\bar{s}}$ (2.9bk)
$\qquad\qquad\qquad\qquad+2\left(\frac{1}{2}s-\frac{1}{4}r\right)\bar{\psi}_{s,\bar{s}}\bar{\phi}_{r-s,-\bar{s}}+2~\bar{\chi}_{s,\bar{s}}\bar{\varphi}_{r-s,-\bar{s}}$
$Q^{(0,\frac{1}{2})}_{r,\bar{r}}=2\pi^{2}i\sum_{s,\bar{s}}\sqrt{2}~\phi_{s,\bar{s}}\chi_{r-s,\bar{r}-\bar{s}}+2\left(\frac{1}{2}\bar{s}-\frac{1}{4}\bar{r}\right)\bar{\chi}_{s,\bar{s}}\bar{\phi}_{r-s,\bar{r}-\bar{s}},$
$\bar{Q}^{(0,\frac{1}{2})}_{\bar{r}}=-2\pi^{2}i\sum_{s,\bar{s}}2\sqrt{2}\left(\frac{1}{2}\bar{s}-\frac{1}{4}\bar{r}\right)\varphi_{s,\bar{s}}\bar{\chi}_{-s,\bar{r}-\bar{s}}+\sqrt{2}~\phi_{s,\bar{s}}\bar{\psi}_{-s,\bar{r}-\bar{s}}$
$\qquad\qquad\qquad\qquad-2\left(\frac{1}{2}\bar{s}-\frac{1}{4}\bar{r}\right)\psi_{s,\bar{s}}\bar{\phi}_{-s,\bar{r}-\bar{s}}-2~\chi_{s,\bar{s}}\bar{\varphi}_{-s,\bar{r}-\bar{s}}$
$\bar{Q}^{(\frac{1}{2},0)}_{r,\bar{r}}=2\pi^{2}i\sum_{s,\bar{s}}\sqrt{2}~\phi_{s,\bar{s}}\bar{\chi}_{r-s,\bar{r}-\bar{s}}-2\left(\frac{1}{2}s-\frac{1}{4}r\right)\chi_{s,\bar{s}}\bar{\phi}_{r-s,\bar{r}-\bar{s}},$
and the $R$ generator is defined as
$R^{(0,0)}_{r,\bar{r}}=\pi^{2}i\sum_{s,\bar{s}}\sqrt{2}~\phi_{s,\bar{s}}\bar{\phi}_{r-s,\bar{r}-\bar{s}}+2~\chi_{s,\bar{s}}\bar{\chi}_{r-s,\bar{r}-\bar{s}}.$ (2.9bl)
The modes transformation is shown in Table 3.3.
Table 6: Field modes transformation under Type I-I BMS4 algebra (2.9ag).
$\bar{\psi}_{s,\bar{s}}$ $\bar{\chi}_{s,\bar{s}}$ $\bar{\varphi}_{s,\bar{s}}$ $\bar{\phi}_{s,\bar{s}}$
$M_{r,\bar{r}}$ $2\left(-\frac{1}{4}\bar{r}-\frac{1}{2}\bar{s}\right)\bar{\chi}_{r+s,\bar{r}+\bar{s}}$ 0 $2\left(-\frac{1}{4}r-\frac{1}{2}s\right)\left(-\frac{1}{4}\bar{r}-\frac{1}{2}\bar{s}\right)\bar{\phi}_{r+s,\bar{r}+\bar{s}}$ 0
$Q^{(\frac{1}{2},0)}_{r}$ $\bar{\varphi}_{r+s,\bar{s}}$ $\left(-\frac{1}{4}r-\frac{1}{2}s\right)\bar{\phi}_{r+s,\bar{s}}$ $\left(-\frac{1}{4}r-\frac{1}{2}s\right)\bar{\psi}_{r+s,\bar{s}}$ $\bar{\chi}_{r+s,\bar{s}}$
$Q^{(0,\frac{1}{2})}_{r,\bar{r}}$ $\left(-\frac{1}{4}\bar{r}-\frac{1}{2}\bar{s}\right)\bar{\phi}_{r+s,\bar{r}+\bar{s}}$ 0 $\left(-\frac{1}{4}\bar{r}-\frac{1}{2}\bar{s}\right)\bar{\chi}_{r+s,\bar{r}+\bar{s}}$ 0
$\bar{Q}^{(0,\frac{1}{2})}_{\bar{r}}$ $\sqrt{2}\left(-\frac{1}{4}\bar{r}-\frac{1}{2}\bar{s}\right)\varphi_{s,\bar{r}+\bar{s}}$ $\frac{1}{\sqrt{2}}\phi_{s,\bar{r}+\bar{s}}$ $\left(-\frac{1}{4}\bar{r}-\frac{1}{2}\bar{s}\right)\psi_{s,\bar{r}+\bar{s}}$ $\chi_{s,\bar{r}+\bar{s}}$
$\bar{Q}^{(\frac{1}{2},0)}_{r,\bar{r}}$ $\frac{1}{\sqrt{2}}\phi_{r+s,\bar{r}+\bar{s}}$ 0 $\left(-\frac{1}{4}r-\frac{1}{2}s\right)\chi_{r+s,\bar{r}+\bar{s}}$ 0
$R^{(0,0)}_{r,\bar{r}}$ $\frac{1}{2}\chi_{r+s,\bar{r}+\bar{s}}$ 0 $\frac{1}{2\sqrt{2}}\phi_{r+s,\bar{r}+\bar{s}}$ 0
The field theory corresponding to these modes is
$\mathcal{L}$ $=\frac{1}{2}\partial_{0}\phi\partial_{0}\phi-\psi_{+}\partial_{0}\psi_{-}-\psi_{-}\partial_{0}\psi_{+}+\psi_{-}\partial_{z}\psi_{-}$ (2.9ca)
$\qquad-\bar{\psi}_{-}\partial_{0}\bar{\psi}_{+}-\bar{\psi}_{+}\partial_{0}\bar{\psi}_{-}+\bar{\psi}_{+}\partial_{\bar{z}}\bar{\psi}_{+}+\bar{\pi}\partial_{0}\bar{\phi}+\partial_{z}\bar{\phi}\partial_{\bar{z}}\bar{\phi}.$
Interestingly, this theory incorporates both the electric Carrollian scalar $\phi$ and the magnetic Carrollian scalars $\bar{\pi}$ and $\bar{\phi}$ [69], providing a total of two bosonic degrees of freedom. The spin and conformal dimensions of these fields are listed as follows
$\phi:l_{\phi}=0,\quad\Delta_{\phi}=\frac{1}{2},$ (2.9cb)
$\psi_{+}:l_{\psi_{+}}=\frac{1}{2},\quad\Delta_{\psi_{+}}=1,$ $\quad\psi_{-}:l_{\psi_{-}}=-\frac{1}{2},\quad\Delta_{\psi_{-}}=1,$
$\bar{\psi}_{+}:l_{\bar{\psi}_{+}}=\frac{1}{2},\quad\Delta_{\bar{\psi}_{+}}=1,$ $\quad\bar{\psi}_{-}:l_{\bar{\psi}_{-}}=-\frac{1}{2},\quad\Delta_{\bar{\psi}_{-}}=1,$
$\bar{\pi}:l_{\bar{\pi}}=0,\quad\Delta_{\bar{\pi}}=\frac{3}{2},$ $\quad\bar{\phi}:l_{\bar{\phi}}=0,\quad\Delta_{\bar{\phi}}=\frac{1}{2},$
The EOMs and mode expansions are respectively
$\partial_{0}^{2}\phi=0,\quad\partial_{0}\psi_{+}=-\partial_{z}\psi_{-},\quad\partial_{0}\psi_{-}=0,$ (2.9cc)
$\partial_{0}\bar{\psi}_{-}=-\partial_{\bar{z}}\bar{\psi}_{+},\quad\partial_{0}\bar{\psi}_{+}=0,\quad\partial_{0}\bar{\pi}=-2\partial_{z}\partial_{\bar{z}}\bar{\phi},\quad\partial_{0}\bar{\phi}=0,$
[…]
$\begin{aligned} &\bar{\pi}(t,z,\bar{z})=\sum_{r,\bar{r}}\bar{\varphi}_{r,\bar{r}}z^{-r-\frac{3}{4}}\bar{z}^{-\bar{r}-\frac{3}{4}}-2\left(r+\frac{1}{4}\right)\left(r+\frac{1}{4}\right)\bar{\phi}_{r,\bar{r}}z^{-r-\frac{1}{4}}\bar{z}^{-\bar{r}-\frac{1}{4}},\\ &\bar{\phi}(t,z,\bar{z})=\sum_{r,\bar{r}}\bar{\phi}_{r,\bar{r}}z^{-r-\frac{1}{4}}\bar{z}^{-\bar{r}-\frac{1}{4}}.\end{aligned}$ (2.9cg)
Realization of magnetic Type II-II algebra (2.9ah)
The modes required to realize the Type II-II algebra are:
$\varphi^{(-\frac{3}{4},-\frac{3}{4})},\quad\phi^{(-\frac{1}{4},-\frac{1}{4})},\quad[\varphi^{(-\frac{3}{4},-\frac{3}{4})}_{r,\bar{r}},\phi^{(-\frac{1}{4},-\frac{1}{4})}_{s,\bar{s}}]=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}},$ (2.9ch)
$\psi^{(-\frac{5}{4},-\frac{3}{4})},\quad\chi^{(\frac{}{4},-\frac{1}{4})},\quad\{\psi^{(-\frac{5}{4},-\frac{3}{4})}_{r,\bar{r}},\chi^{(\frac{1}{4},-\frac{1}{4})}_{s,\bar{s}}\}=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}},$
$\bar{\psi}^{(-\frac{3}{4},-\frac{5}{4})},\quad\bar{\chi}^{(-\frac{1}{4},\frac{1}{4})},\quad\{\bar{\psi}^{(-\frac{3}{4},-\frac{5}{4})}_{r,\bar{r}},\bar{\chi}^{(-\frac{1}{4},\frac{1}{4})}_{s,\bar{s}}\}=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}},$
$\bar{\varphi}^{(-\frac{5}{4},-\frac{5}{4})},\quad\bar{\phi}^{(\frac{1}{4},\frac{1}{4})},\quad[\bar{\varphi}^{(-\frac{5}{4},-\frac{5}{4})}_{r,\bar{r}},\bar{\phi}^{(\frac{1}{4},\frac{1}{4})}_{s,\bar{s}}]=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}}.$
The algebra is constructed as
$L_{m}=-4\pi^{2}i\sum_{s,\bar{s}}\left(s-\frac{1}{4}m\right)\varphi_{s,\bar{s}}\phi_{m-s,-\bar{s}}-\left(s+\frac{1}{4}m\right)\psi_{s,\bar{s}}\chi_{m-s,-\bar{s}}$ (2.9ci)
$\qquad\qquad\qquad\qquad-\left(s-\frac{1}{4}m\right)\bar{\psi}_{s,\bar{s}}\bar{\chi}_{m-s,-\bar{s}}+\left(s+\frac{1}{4}m\right)\bar{\varphi}_{s,\bar{s}}\bar{\phi}_{m-s,-\bar{s}},$
$\bar{L}_{\bar{m}}=-4\pi^{2}i\sum_{s,\bar{s}}\left(\bar{s}-\frac{1}{4}\bar{m}\right)\varphi_{s,\bar{s}}\phi_{-s,\bar{m}-\bar{s}}-\left(\bar{s}-\frac{1}{4}\bar{m}\right)\psi_{s,\bar{s}}\chi_{-s,\bar{m}-\bar{s}}$
$\qquad\qquad\qquad\qquad-\left(\bar{s}+\frac{1}{4}\bar{m}\right)\bar{\psi}_{s,\bar{s}}\bar{\chi}_{-s,\bar{m}-\bar{s}}+\left(\bar{s}+\frac{1}{4}\bar{m}\right)\bar{\varphi}_{s,\bar{s}}\bar{\phi}_{-s,\bar{m}-\bar{s}},$
$M_{r,\bar{r}}=2\pi^{2}i\sum_{s,\bar{s}}\phi_{s,\bar{s}}\phi_{r-s,\bar{r}-\bar{s}},$
$Q^{(\frac{1}{2},0)}_{r}=-4\pi^{2}i\sum_{s,\bar{s}}\varphi_{s,\bar{s}}\chi_{r-s,-\bar{s}}-\left(\frac{1}{2}s-\frac{3}{4}r\right)\phi_{s,\bar{s}}\psi_{r-s,-\bar{s}}$ (2.9cj)
$\qquad\qquad\qquad\qquad+\bar{\psi}_{s,\bar{s}}\bar{\phi}_{r-s,-\bar{s}}-\left(\frac{1}{2}s-\frac{3}{4}r\right)\bar{\varphi}_{r-s,-\bar{s}}$
$Q^{(1,\frac{1}{2})}_{r,\bar{r}}=-4\pi^{2}i\sum_{s,\bar{s}}\phi_{s,\bar{s}}\chi_{r-s,\bar{r}-\bar{s}},$
$\bar{Q}^{(0,\frac{1}{2})}_{\bar{r}}=-4\pi^{2}i\sum_{s,\bar{s}}\varphi_{s,\bar{s}}\bar{\chi}_{-s,\bar{r}-\bar{s}}-\left(\frac{1}{2}\bar{s}-\frac{3}{4}\bar{r}\right)\phi_{s,\bar{s}}\bar{\psi}_{-s,\bar{r}-\bar{s}}$
$\qquad\qquad\qquad\qquad-\psi_{s,\bar{s}}\bar{\phi}_{-s,\bar{r}-\bar{s}}+\left(\frac{1}{2}\bar{s}-\frac{3}{4}\bar{r}\right)\chi_{s,\bar{s}}\bar{\varphi}_{-s,\bar{r}-\bar{s}}$
$\bar{Q}^{(\frac{1}{2},1)}_{r,\bar{r}}=2\pi^{2}i\sum_{s,\bar{s}}\sqrt{2}~\phi_{s,\bar{s}}\bar{\chi}_{r-s,\bar{r}-\bar{s}},$
and
$R^{(1,1)}_{r,\bar{r}}=-4\pi^{2}i\sum_{s,\bar{s}}\phi_{s,\bar{s}}\bar{\phi}_{r-s,\bar{r}-\bar{s}}+\chi_{s,\bar{s}}\bar{\chi}_{r-s,\bar{r}-\bar{s}}.$ (2.9ck)
The transformation of these modes under the algebra is summarized in Table 3.3.
Table 7: Field modes transformation under Type II-II BMS4 algebra (2.9ah).
$\bar{\psi}_{s,\bar{s}}$ $\bar{\chi}_{s,\bar{s}}$ $\bar{\varphi}_{s,\bar{s}}$ $\bar{\phi}_{s,\bar{s}}$
$M_{r,\bar{r}}$ 0 0 0 0
$Q^{(\frac{1}{2},0)}_{r}$ $\left(-\frac{3}{4}r-\frac{1}{2}s\right)\bar{\varphi}_{r+s,\bar{s}}$ $\bar{\phi}_{r+s,\bar{s}}$ $\bar{\psi}_{r+s,\bar{s}}$ $\left(\frac{1}{4}r-\frac{1}{2}s\right)\bar{\chi}_{r+s,\bar{s}}$
$Q^{(1,\frac{1}{2})}_{r,\bar{r}}$ 0 0 0 0
$\bar{Q}^{(0,\frac{1}{2})}_{\bar{r}}$ $\varphi_{s,\bar{r}+\bar{s}}$ $\left(\frac{1}{4}\bar{r}-\frac{1}{2}\bar{s}\right)\phi_{s,\bar{r}+\bar{s}}$ $\psi_{s,\bar{r}+\bar{s}}$ $\left(\frac{1}{4}\bar{r}-\frac{1}{2}\bar{s}\right)\chi_{s,\bar{r}+\bar{s}}$
$\bar{Q}^{(\frac{1}{2},1)}_{r,\bar{r}}$ $\phi_{r+s,\bar{r}+\bar{s}}$ 0 0 0
$R^{(1,1)}_{r,\bar{r}}$ $\chi_{r+s,\bar{r}+\bar{s}}$ 0 $\phi_{r+s,\bar{r}+\bar{s}}$ 0
The constructed theory is
$\mathcal{L}$ $=\frac{1}{2}\partial_{0}\phi\partial_{0}\phi-\psi_{+}\partial_{0}\psi_{-}-\psi_{-}\partial_{0}\psi_{+}-\bar{\psi}_{-}\partial_{0}\bar{\psi}_{+}-\bar{\psi}_{+}\partial_{0}\bar{\psi}_{-}+\bar{\varphi}\partial_{0}\bar{\phi}.$ (2.9cz)
The spin and conformal dimension of the fields are given by
$\phi:l_{\phi}=0,\quad\Delta_{\phi}=\frac{1}{2},$ (2.9da)
$\psi_{+}:l_{\psi_{+}}=\frac{1}{2},\quad\Delta_{\psi_{+}}=2,$ $\quad\psi_{-}:l_{\psi_{-}}=-\frac{1}{2},\quad\Delta_{\psi_{-}}=0,$
$\bar{\psi}_{+}:l_{\bar{\psi}_{+}}=\frac{1}{2},\quad\Delta_{\bar{\psi}_{+}}=0,$ $\quad\bar{\psi}_{-}:l_{\bar{\psi}_{-}}=-\frac{1}{2},\quad\Delta_{\bar{\psi}_{-}}=2,$
$\bar{\varphi}:l_{\bar{\varphi}}=0,\quad\Delta_{\bar{\varphi}}=-\frac{1}{2},$ $\quad\bar{\phi}:l_{\bar{\phi}}=0,\quad\Delta_{\bar{\phi}}=\frac{5}{2},$
The EOMs and mode expansions are respectively
$\partial_{0}^{2}\phi=0,\quad\partial_{0}\psi_{+}=\partial_{0}\psi_{-}=0,\quad\partial_{0}\bar{\psi}_{+}=\partial_{0}\bar{\psi}_{-}=0,\quad\partial_{0}\bar{\varphi}=\partial_{0}\bar{\phi}=0,$ (2.9db)
[…]
$\begin{aligned} &\bar{\varphi}(t,z,\bar{z})=\sum_{r,\bar{r}}\bar{\varphi}_{r,\bar{r}}z^{-r+\frac{1}{4}}\bar{z}^{-\bar{r}+\frac{1}{4}},\quad\bar{\phi}(t,z,\bar{z})=\sum_{r,\bar{r}}\bar{\phi}_{r,\bar{r}}z^{-r-\frac{5}{4}}\bar{z}^{-\bar{r}-\frac{5}{4}}.\end{aligned}$ (2.9df)
Realization of magnetic Type I-II algebra (2.9ai)
Finally, we construct the theory exhibiting Type I-II symmetry. The corresponding modes are given by:
$\varphi^{(-\frac{3}{4},-\frac{3}{4})},\quad\phi^{(-\frac{1}{4},-\frac{1}{4})},\quad[\varphi^{(-\frac{3}{4},-\frac{3}{4})}_{r,\bar{r}},\phi^{(-\frac{1}{4},-\frac{1}{4})}_{s,\bar{s}}]=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}},$ (2.9dg)
$\psi^{(-\frac{1}{4},-\frac{3}{4})},\quad\chi^{(-\frac{3}{4},-\frac{1}{4})},\quad\{\psi^{(-\frac{1}{4},-\frac{3}{4})}_{r,\bar{r}},\chi^{(-\frac{3}{4},-\frac{1}{4})}_{s,\bar{s}}\}=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}},$
$\bar{\psi}^{(-\frac{3}{4},-\frac{5}{4})},\quad\bar{\chi}^{(-\frac{1}{4},\frac{1}{4})},\quad\{\bar{\psi}^{(-\frac{3}{4},-\frac{5}{4})}_{r,\bar{r}},\bar{\chi}^{(-\frac{1}{4},\frac{1}{4})}_{s,\bar{s}}\}=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}},$
$\bar{\varphi}^{(-\frac{1}{4},-\frac{5}{4})},\quad\bar{\phi}^{(-\frac{3}{4},\frac{1}{4})},\quad[\bar{\varphi}^{(-\frac{1}{4},-\frac{5}{4})}_{r,\bar{r}},\bar{\phi}^{(-\frac{3}{4},\frac{1}{4})}_{s,\bar{s}}]=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}}.$
The algebra is constructed as:
$L_{m}=-4\pi^{2}i\sum_{s,\bar{s}}\left(s-\frac{1}{4}m\right)\varphi_{s,\bar{s}}\phi_{m-s,-\bar{s}}-\left(s-\frac{3}{4}m\right)\psi_{s,\bar{s}}\chi_{m-s,-\bar{s}}$ (2.9dh)
$\qquad\qquad\qquad\qquad-\left(s-\frac{1}{4}m\right)\bar{\psi}_{s,\bar{s}}\bar{\chi}_{m-s,-\bar{s}}+\left(s-\frac{3}{4}m\right)\bar{\varphi}_{s,\bar{s}}\bar{\phi}_{m-s,-\bar{s}},$
$\bar{L}_{\bar{m}}=-4\pi^{2}i\sum_{s,\bar{s}}\left(\bar{s}-\frac{1}{4}\bar{m}\right)\varphi_{s,\bar{s}}\phi_{-s,\bar{m}-\bar{s}}-\left(\bar{s}-\frac{1}{4}\bar{m}\right)\psi_{s,\bar{s}}\chi_{-s,\bar{m}-\bar{s}}$
$\qquad\qquad\qquad\qquad-\left(\bar{s}+\frac{1}{4}\bar{m}\right)\bar{\psi}_{s,\bar{s}}\bar{\chi}_{-s,\bar{m}-\bar{s}}+\left(\bar{s}+\frac{1}{4}\bar{m}\right)\bar{\varphi}_{s,\bar{s}}\bar{\phi}_{-s,\bar{m}-\bar{s}},$
$M_{r,\bar{r}}=2\pi^{2}i\sum_{s,\bar{s}}\phi_{s,\bar{s}}\phi_{r-s,\bar{r}-\bar{s}}+2\left(\frac{1}{2}\bar{s}-\frac{1}{4}\bar{m}\right)\chi_{s,\bar{s}}\chi_{r-s,\bar{m}-\bar{s}},$
the fermionic generators are
$Q^{(\frac{1}{2},0)}_{r}=-2\sqrt{2}\pi^{2}i\sum_{s,\bar{s}}2\left(\frac{1}{2}s-\frac{1}{4}r\right)\varphi_{s,\bar{s}}\chi_{r-s,-\bar{s}}+\phi_{s,\bar{s}}\psi_{r-s,-\bar{s}}$ (2.9di)
$\qquad\qquad\qquad\qquad-2\left(\frac{1}{2}s-\frac{1}{4}r\right)\bar{\psi}_{s,\bar{s}}\bar{\phi}_{r-s,-\bar{s}}-\bar{\chi}_{s,\bar{s}}\bar{\varphi}_{r-s,-\bar{s}}$
$Q^{(0,\frac{1}{2})}_{r,\bar{r}}=2\sqrt{2}\pi^{2}i\sum_{s,\bar{s}}\phi_{s,\bar{s}}\chi_{r-s,\bar{r}-\bar{s}},$
$\bar{Q}^{(0,\frac{1}{2})}_{\bar{r}}=-4\pi^{2}i\sum_{s,\bar{s}}\varphi_{s,\bar{s}}\bar{\chi}_{-s,\bar{r}-\bar{s}}-\left(\frac{1}{2}\bar{s}-\frac{3}{4}\bar{r}\right)\phi_{s,\bar{s}}\bar{\psi}_{-s,\bar{r}-\bar{s}}$
$\qquad\qquad\qquad\qquad+\psi_{s,\bar{s}}\bar{\phi}_{-s,\bar{r}-\bar{s}}-\left(\frac{1}{2}\bar{s}-\frac{3}{4}\bar{r}\right)\chi_{s,\bar{s}}\bar{\varphi}_{-s,\bar{r}-\bar{s}}$
$\bar{Q}^{(\frac{1}{2},1)}_{r,\bar{r}}=-4\pi^{2}i\sum_{s,\bar{s}}\phi_{s,\bar{s}}\bar{\chi}_{r-s,\bar{r}-\bar{s}}+2\left(\frac{1}{2}s-\frac{1}{4}r\right)\chi_{s,\bar{s}}\bar{\phi}_{r-s,\bar{r}-\bar{s}},$
and the bosonic $R$ generators are
$R^{(0,1)}_{r,\bar{r}}=2\sqrt{2}\pi^{2}i\sum_{s,\bar{s}}\phi_{s,\bar{s}}\bar{\phi}_{r-s,\bar{r}-\bar{s}}-\chi_{s,\bar{s}}\bar{\chi}_{r-s,\bar{r}-\bar{s}}.$ (2.9dj)
The modes transformation is shown in Table 3.3.
Table 8: Field modes transformation under Type I-II BMS4 algebra (2.9ai).
$\bar{\psi}_{s,\bar{s}}$ $\bar{\chi}_{s,\bar{s}}$ $\bar{\varphi}_{s,\bar{s}}$ $\bar{\phi}_{s,\bar{s}}$
$M_{r,\bar{r}}$ 0 0 0 0
$Q^{(\frac{1}{2},0)}_{r}$ $\frac{1}{\sqrt{2}}\bar{\varphi}_{r+s,\bar{s}}$ $\sqrt{2}\left(-\frac{1}{4}r-\frac{1}{2}s\right)\bar{\phi}_{r+s,\bar{s}}$ $\sqrt{2}\left(-\frac{1}{4}r-\frac{1}{2}s\right)\bar{\psi}_{r+s,\bar{s}}$ $\frac{1}{\sqrt{2}}\bar{\chi}_{r+s,\bar{s}}$
$Q^{(0,\frac{1}{2})}_{r,\bar{r}}$ 0 0 0 0
$\bar{Q}^{(0,\frac{1}{2})}_{\bar{r}}$ $\varphi_{s,\bar{r}+\bar{s}}$ $\left(\frac{1}{4}\bar{r}-\frac{1}{2}\bar{s}\right)\phi_{s,\bar{r}+\bar{s}}$ $\psi_{s,\bar{r}+\bar{s}}$ $\left(\frac{1}{4}\bar{r}-\frac{1}{2}\bar{s}\right)\chi_{s,\bar{r}+\bar{s}}$
$\bar{Q}^{(\frac{1}{2},1)}_{r,\bar{r}}$ $\phi_{r+s,\bar{r}+\bar{s}}$ 0 $\left(-\frac{1}{4}r-\frac{1}{2}s\right)\chi_{r+s,\bar{r}+\bar{s}}$ 0
$R^{(0,1)}_{r,\bar{r}}$ $\frac{1}{\sqrt{2}}\chi_{r+s,\bar{r}+\bar{s}}$ 0 $\frac{1}{\sqrt{2}}\phi_{r+s,\bar{r}+\bar{s}}$ 0
The field theory incorporating these modes is
$\mathcal{L}$ $=\frac{1}{2}\partial_{0}\phi\partial_{0}\phi-\psi_{+}\partial_{0}\psi_{-}-\psi_{-}\partial_{0}\psi_{+}+\psi_{-}\partial_{z}\psi_{-}-\bar{\psi}_{-}\partial_{0}\bar{\psi}_{+}-\bar{\psi}_{+}\partial_{0}\bar{\psi}_{-}+A_{+}\partial_{0}A_{-}.$ (2.9dy)
It is worth noting that this nonchiral symmetric theory includes spin-$\pm 1$ bosons $A_{\pm}$. Especially, the $A_{\pm}$ $\bar{\psi}_{\pm}$ part of this theory is single copy of the theory studied in [21] as perturbative supergravity. The spin and conformal dimensions of the fields are summarized as
$\phi:l_{\phi}=0,\quad\Delta_{\phi}=\frac{1}{2},$ (2.9dz)
$\psi_{+}:l_{\psi_{+}}=\frac{1}{2},\quad\Delta_{\psi_{+}}=1,$ $\quad\psi_{-}:l_{\psi_{-}}=-\frac{1}{2},\quad\Delta_{\psi_{-}}=1,$
$\bar{\psi}_{+}:l_{\bar{\psi}_{+}}=\frac{1}{2},\quad\Delta_{\bar{\psi}_{+}}=0,$ $\quad\bar{\psi}_{-}:l_{\bar{\psi}_{-}}=-\frac{1}{2},\quad\Delta_{\bar{\psi}_{-}}=2,$
$A_{+}:l_{A_{+}}=1,\quad\Delta_{A_{+}}=\frac{1}{2},$ $\quad A_{-}:l_{A_{-}}=-1,\quad\Delta_{A_{-}}=\frac{3}{2},$
The EOMs and mode expansions are respectively
$\partial_{0}^{2}\phi=0,\quad\partial_{0}\psi_{+}=-\partial_{z}\psi_{-},\quad\partial_{0}\psi_{-}=0,$ (2.9ea)
$\partial_{0}\bar{\psi}_{-}=\partial_{0}\bar{\psi}_{+}=0,\quad\partial_{0}A_{+}=\partial_{0}A_{-}=0,$
[…]
$\begin{aligned} &A_{+}(t,z,\bar{z})=\sum_{r,\bar{r}}\bar{\varphi}_{r,\bar{r}}z^{-r-\frac{3}{4}}\bar{z}^{-\bar{r}+\frac{1}{4}},\quad A_{-}(t,z,\bar{z})=\sum_{r,\bar{r}}\bar{\phi}_{r,\bar{r}}z^{-r-\frac{1}{4}}\bar{z}^{-\bar{r}-\frac{5}{4}}.\end{aligned}$ (2.9ee)
In the preceding constructions, we have realized the magnetic super BMS4 algebras through the theories (2.9at), (2.9be), (2.9ca), (2.9cz), and (2.9dy). Notably, there is no single simple theory that accommodates all types of magnetic algebras, highlighting their nontrivial structure compared to the electric algebras. Theories (2.9ca), (2.9cz), and (2.9dy) are not standard extended supersymmetric theories; nevertheless, the inclusion of $R$-symmetry is essential for the closure of the algebra. From this perspective, the left-hand and right-hand fermionic generators can be viewed as independent symmetry algebras, and their combinations should be regarded as extensions of the superalgebra. In each case, the $R$-symmetry maps bosons to bosons and fermions to fermions. It is also noteworthy that in (2.9ca), the presence of the Carrollian magnetic scalar demonstrates that, for supersymmetric BMS4 theories, the electric and magnetic sectors are equally significant and are related via $R$-symmetry. In the case of (2.9dy), the emergence of spin-$1$ fields corresponds to massless vector fields in the bulk 44D theory, rendering this nontrivial theory particularly relevant for future discussions in flat holography.
4 Discussion
In this paper, we conducted a systematic study of the supersymmetric extensions of the BMS4 algebra. In Section 2, we explored all possible electric and magnetic super BMS4 algebras that admit finite-dimensional subalgebras. For the electric case, we imposed the condition that the anticommutator of supercharges contains only supertranslations $M$, whereas for the magnetic case, the inclusion of superrotations $L$ and $\bar{L}$ is required. Our analysis revealed ten distinct electric super BMS4 algebras, two types of chiral magnetic super BMS4 algebras, and four nonchiral magnetic superalgebras formed from combinations of two chiral superalgebras.
In Section 3, we implemented an inverse construction to realize these superalgebras in free field theories. Notably, we identified a simple theory (2.9y) that realizes (2.9s), (2.9t), and (2.9u), consisting of one electric Carrollian scalar and two electric Carrollian spinors. For magnetic supersymmetry, there is a unique realization for each type. Although the of theory (2.9be) resembles the electric theory (2.9y), the conformal dimensions of the fermions differ, highlighting the distinct nature of magnetic constructions. The significance of magnetic Carrollian scalars is particularly evident in theory (2.9ca), while theory (2.9dy) plays a key role in the study of theories with bulk electromagnetic fields. Especially, part of (2.9dy) was studied in [21] as model of perturbative supergravity. The constructions presented in Section 3 focus on the simplest scalar supermultiplets, but the framework allows straightforward study of more general multiplets for further applications.
The global part of the BMS4 algebra is isomorphic to both the 44D Poincaré algebra and the 33D Carrollian conformal algebra. Similar relations hold for their supersymmetric extensions and the corresponding field theory realizations. As a supersymmetric extension of the 44D Poincaré algebra, it is imposed strict constraints by Haag–Łopuszański–Sohnius theorem. Indeed, the super Poincaré algebra coincides with the global part of the electric super BMS4 algebra (2.9s). The remaining electric superalgebras discussed in Section 2.2 involve supercharges with spin greater than $\tfrac{1}{2}$, and therefore their global parts do not correspond to physically realized superalgebras. Similarly, the magnetic superalgebras, whose supercharge anticommutators also produce superrotations, lack a direct physical superalgebra interpretation for their global sectors. Nevertheless, this does not diminish their relevance in the context of flat holography. On the contrary, these algebras highlight how supersymmetric extensions of BMS4 can go beyond the strict confines of super Poincaré, potentially capturing structures that only emerge in asymptotic or holographic regimes. In particular, such algebras may arise from large-radius limits of extended super asymptotic symmetries, and the appearance of $R$-symmetry offers a novel perspective on how internal symmetries interplay with asymptotic supersymmetry.
The BMS4 algebra is closely related to the 33D Carrollian conformal algebra in both algebraic and field-theoretic aspects. By requiring that supercharges admit finite-dimensional subsectors, the global part of all superalgebras discussed in this work can be naturally reinterpreted as 33D super Carrollian conformal algebras. This dual perspective not only unifies the classification but also connects the free field realizations to a broader Carrollian framework. Indeed, all the free theories constructed here can be viewed as super Carrollian conformal theories containing fields of diverse spins and conformal dimensions. The interplay between electric and magnetic Carrollian scalar theories, related by the $R$-symmetry of the Type I-I magnetic superalgebra, further illustrates how Carrollian holography accommodates supersymmetric structures beyond conventional relativistic situations. In this way, our results underline the role of Carrollian methods as a powerful organizing language for flat space holography and pave the way for future explorations of supersymmetric Carrollian conformal field theories [70].
There are several promising directions for extending this study:
Super BMS4 algebra from asymptotic symmetry. The origin of the BMS4 algebra lies in the study of the symmetries of asymptotically flat gravity. In a similar manner, the super BMS4 algebra can also be derived from this perspective [71, 72, 73, 74, 40]. It would be interesting to extend such an asymptotic analysis to the new super BMS4 algebras discussed in this work, and further investigate the physical interpretation of these new superalgebras.
Extended supersymmetry. Extended super BMS3 and BMS4 algebras with $\mathcal{N}$ copy of electric super charges was well studied in the literature [75, 76, 73, 74, 77, 78, 51, 55]. The presence of $R$-symmetry in the magnetic Type I-I, Type II-II, and Type I-II algebras indicates that these should also be regarded as extended superalgebras. Interestingly, in general superalgebras, the $R$-symmetry can carry nonzero spin. It is straightforward to consider the case of conventional extended supersymmetry, i.e., multiple sets of the same supercharges, including both electric and magnetic types. In such scenarios, $R$-symmetries with generic spin are naturally expected in general extended super BMS4 algebras. Furthermore, the physical interpretation of these $R$-symmetries, particularly those with spin, is an intriguing question. From a holographic perspective, such novel $R$-symmetries may correspond to symmetries associated with compactified dimensions, thereby providing a bridge between asymptotic flat space physics and string/M-theory constructions.
More general supersymmetry. In this work, we focused on supercharges in the charge representation (2.3). A more general supersymmetric extension may admit charges in arbitrary representations, including structures reminiscent of $w$-algebras [48, 79]. It is expected that super BMS4 algebras with higher-spin generators are naturally accommodated within a $w$-algebra framework, which could have direct applications in the study of soft theorems and memory effects. Extending the present framework to encompass higher-spin symmetries would not only generalize the classification but also provide algebraic tools for organizing scattering data in flat space, in close analogy with the role of $W$-symmetry in 22D CFT.
Holography with supersymmetry. This paper has classified all possible super BMS4 algebras in the charge representation with finite subalgebras, providing a firm foundation for understanding supersymmetry in flat holography. These results serve as fundamental building blocks for constructing Fock spaces, boundary field theories, and analyzing constraints imposed by supersymmetry in holographic setups. In particular, the structure of $R$-symmetry offers guidance for constructing full 1010D spaces of string theories in flat holography, potentially shedding light on the embedding of flat space holography into string theory. Additionally, the results presented here suggest new directions for celestial holography: free superfield theories allow for explicit computation of correlation functions, thereby establishing bridges between supersymmetric scattering amplitudes and celestial correlators. Notably, the theory (2.9dy), which contains a massless spin-$1$ particle, is directly relevant for the study of soft photon theorems and may provide a testing ground for exploring how supersymmetry constrains celestial amplitudes.
Acknowledgments
We are deeply grateful to Bin Chen and Sergio Cecotti for valuable discussions.
Appendix A A Tensor Product of Virasoro Algebra Representations
In this appendix we show the solutions of equation (2.8) and (2.9e).
For convenience, we repeat equation (2.8) here:
$(l_{k}n-(r_{i}+r_{j}))f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})-(l_{j}n-r_{j})f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j}+n)-(l_{i}n-r_{i})f_{l_{i},l_{j}}^{l_{k}}(r_{i}+n,r_{j})=0.$ (2.9a)
This is a difference equation, and solving it analytically for generic values of the parameters $l_{i}$, $l_{j}$, and $l_{k}$ is challenging. Instead, we proceed by obtaining numerical solutions for specific parameter choices. The setup is as follows:
[…]
A solution is referred as continuous if $f$ is a continuous function, and denoted as discrete if $f$ involves Kronecker $\delta$ function. The results are summarized as follows
continuous
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=1,$ $l_{k}=l_{i}+l_{j},$ (2.9ca)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=l_{j}r_{i}-l_{i}r_{j},$ $l_{k}=l_{i}+l_{j}-1,$ (2.9cb)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=r_{i}(l_{j}r_{i}+r_{j}),$ $l_{i}=0,\quad l_{k}=l_{j}-2,$ (2.9cc)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=(r_{i}+r_{j})(l_{j}r_{i}-l_{i}r_{j}),$ $l_{i}+l_{j}=1,\quad l_{k}=-1,$ (2.9cd)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=r_{i},$ $l_{i}=l_{j}=0,\quad l_{k}=-1,$ (2.9ce)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=l_{j}r_{i}-l_{i}r_{j},$ $l_{i}=l_{j}=0,\quad l_{k}=-3,$ (2.9cf)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=r_{i}(r_{i}+r_{j})(2r_{i}+r_{j}),$ $l_{i}=0,\quad l_{j}=2,\quad l_{k}=-1.$ (2.9cg)
discrete
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=\delta_{r_{i}+r_{j}},$ $l_{i}+l_{j}=-1,\quad l_{k}=0,$ (2.9da)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=\delta_{r_{i}},$ $l_{i}=-1,\quad l_{j}=l_{k},$ (2.9db)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=r_{i}\delta(r_{i}+r_{j}),$ $l_{i}=l_{j}=l_{k}=0,$ (2.9dc)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=r_{i}\delta_{r_{j}},$ $l_{i}=0,\quad l_{k}=l_{j}=-1,$ (2.9dd)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=\delta_{r_{i}}\delta_{r_{j}},$ $l_{i}=l_{j}=-1,\quad l_{k}=0,$ (2.9de)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=\frac{1}{r_{i}}\delta_{r_{i}+r_{j}}+\frac{1}{r_{j}}\delta_{r_{i}}-\frac{1}{r_{i}}\delta_{r_{j}},$ $l_{i}=l_{j}=-1,\quad l_{k}=0.$ (2.9df)
These solutions remain valid under the exchange of $i$ and $j$. Among them, (2.9ca), (2.9cb), and (2.9cc) coincide with the solutions obtained in (2.9). Solution (2.9df) does not play a role in the construction of the superalgebras. And the other solutions are only consistent for negative parameter values, so they are not in our interest for constructing the superalgebras.
It is worth emphasizing that (2.9da) is the solution used in (2.9h), which automatically satisfied by the canonical commutation relations of the field modes. For instance, (2.9i) requires $l_{\varphi}+l_{\phi}=-1$ and $\bar{l}_{\varphi}+\bar{l}_{\phi}=-1$, while the spin indices of the identity operator are $l_{I}=\bar{l}_{I}=0$.
The difference equation for the field modes (2.9d) is repeated here:
$((l_{i}+1)m-r)c(r-m,t)+(l_{j}m-t+r)c(r,t)-(l_{Q}m-t)c(r,t+m)=0.$ (2.9e)
The numerical method we used are similar, but the setup is
[…]
Using this setup, we find the following solutions
continuous
$c_{l_{i},l_{j}}^{l_{k}}(r,n)=1,$ $l_{k}=l_{i}+l_{j}+1,$ (2.9ga)
$c_{l_{i},l_{j}}^{l_{k}}(r,n)=l_{k}r-(l_{i}+1)n,$ $l_{k}=l_{i}+l_{j}+2,$ (2.9gb)
$c_{l_{i},l_{j}}^{l_{k}}(r,n)=r((2+l_{j})r-n),$ $l_{i}=-1,\quad l_{k}=l_{j}+2,$ (2.9gc)
$c_{l_{i},l_{j}}^{l_{k}}(r,n)=n(r+(l_{i}+1)n),$ $l_{i}+l_{j}=-3,\quad l_{k}=0,$ (2.9gd)
discrete
$c_{l_{i},l_{j}}^{l_{k}}(r,n)=\delta_{r},$ $l_{i}=0,\quad l_{k}=l_{j},$ (2.9ha)
$c_{l_{i},l_{j}}^{l_{k}}(r,n)=n\delta_{r},$ $l_{i}=l_{k}=0,\quad l_{j}=-1,$ (2.9hb)
$c_{l_{i},l_{j}}^{l_{k}}(r,n)=\frac{1}{n}\delta_{n-r}-\frac{1}{r}\delta_{n}-\frac{1}{n}\delta_{r},$ $l_{i}=l_{j}=0,\quad l_{k}=-1,$ (2.9hc)
$c_{l_{i},l_{j}}^{l_{k}}(r,n)=\delta_{r}\delta_{n},$ $l_{i}=l_{j}=0,\quad l_{k}=-1,$ (2.9hd)
$c_{l_{i},l_{j}}^{l_{k}}(r,n)=\delta_{n},$ $l_{i}+l_{j}=l_{k}=-1,$ (2.9he)
$c_{l_{i},l_{j}}^{l_{k}}(r,n)=r\delta_{n},$ $l_{i}=l_{j}=l_{k}=-1.$ (2.9hf)
These solutions remain valid under the exchange $l_{i}\leftrightarrow l_{j}$ accompanied by the shift $n\to n-r$. Among them, only (2.9ga) and (2.9gb), which correspond to (2.9e), are relevant for constructing the field modes.